### v1

# On Fourier Asymptotics and Effective Equidistribution

Abstract.
We prove effective equidistribution of expanding horocycles in $\mathrm{SL_{2}(\mathbb{Z})}\backslash\mathrm{SL_{2}(\mathbb{R})}$ with respect to various classes of Borel probability measures on $\mathbb{R}$ having certain Fourier asymptotics. Our proof involves new techniques combining tools from automorphic forms and harmonic analysis.
In particular, we prove effective equidistribution with respect to the Hausdorff measures on certain shifted, possibly irrational, missing digit Cantor sets, whose Hausdorff dimensions can be as small as 0.61. This improves upon the dimension and rationality restrictions of the recent breakthrough of Khalil–Luethi [Invent. Math, 232(2): 713–831, 2023].
As an application in Diophantine approximation, a complete Khintchine’s theorem follows for a large class of self-similar measures extending a groundbreaking work of Yu [arXiv, 2101.05910].
1. Introduction
1.1. Equidistribution of horocycle flow
Dynamics on $\mathbb{X}:=\operatorname{SL}_{2}(\mathbb{Z})\backslash\operatorname{SL}_{2}(\mathbb{R})$, the unit cotangent bundle of the modular surface, is a central topic lying in the interface of homogeneous dynamics and representation theory and has numerous applications in various number theoretic problems. We consider the homogeneous space $\mathbb{X}$ equipped with the $\operatorname{SL}_{2}(\mathbb{R})$-invariant probability measure $m_{\mathbb{X}}$. Let $g_{y}:=\left(\begin{smallmatrix}\sqrt{y}&\\ &\sqrt{y^{-1}}\end{smallmatrix}\right)$ for $y>0$ be the geodesic flow and the corresponding expanding horocycle flow is given by $n(x):=\bigl{(}\begin{smallmatrix}1&x\\ 0&1\end{smallmatrix}\bigr{)}$ for $x\in\mathbb{R}$. Let $\mu$ denote a Borel probability measure on $\mathbb{R}$. Let $\mu_{y}:=\int\delta_{n(x)g_{y}}\,\mathrm{d}\mu(x)$ denote the probability measure on $\mathbb{X}$ supported on the horocycle of length of $y^{-1}$, that is, for $\phi\in C_{c}^{\infty}(\mathbb{X})$ we write
$\mu_{y}(\phi):=\int\phi(n(x)g_{y})\,\mathrm{d}\mu(x).$ (1.1)
We record the following (folklore) problem which is currently mostly wide open.
Problem 1.
Given a Borel probability measure $\mu$ on $\mathbb{R}$, determine whether $\mu_{y}\xrightarrow{\text{weak-}^{\ast}}m_{\mathbb{X}}$ as $y\to 0$ with a polynomial order convergence rate, that is, if there is an $\eta>0$ such that
$\mu_{y}(\phi)=m_{\mathbb{X}}(\phi)+O_{\mu,\phi}(y^{\eta}),\quad\phi\in C_{c}^{\infty}(\mathbb{X}),$
as $y\to 0$ where $m_{\mathbb{X}}(\phi):=\int\phi\,\mathrm{d}m_{\mathbb{X}}$.
Let $\widehat{\mu}$ denote the Fourier transform of $\mu$; see(2.4). The goal of this paper is to initiate a study of Problem 1, depending on the asymptotic behavior of $\widehat{\mu}$. Loosely speaking, we answer Problem 1 affirmatively for $\mu$ with $\widehat{\mu}$ having
  1. (1)
    no pointwise decay, but a certain polynomial decay on average,
  2. (2)
    arbitrarily slow polynomial decay, but having precise oscillatory asymptotics.
Interestingly, both of the above types contain many interesting class of measures. Recently, in [12] and [58], a large class of self-similar measures obtained that fall under kind 1 as above. On the other hand, the push-forward of the Lebesgue measure by analytic non-constant maps fall under kind 2 as above.
Problem 1 has a long history when $\mu$ is the normalized Lebesgue measure supported on $[0,1]$. In this case, Problem 1 has been investigated by numerous people over the last decades, starting with Selberg and Zagier; see [44] and [52, 51, 18, 36, 11, 45] that make the equidistribution result by Dani–Smillie [15] (and also [19]) effective.
For absolutely continuous measures, it is known that equidistribution holds as in Problem 1, but nothing can be said, a priori, about the rate of convergence. In fact, the convergence could happen arbitrary slowly if the density function (coming from Radon–Nikodym theorem) does not have enough regularity. Kleinbock–Margulis, in [32, Theorem 1.1], answered Problem 1 affirmatively for absolutely continuous measures with smooth compactly supported densities in the space of matrices. Recently, Björklund–Gorodnik in [7] extended [32] to continuous densities, among other things.
Problem 1 is significantly difficult when the measure is not absolutely continuous. Such problem emerges from deep works by Kleinbock–Margulis [31] and Kleinbock–Lindenstrauss–Weiss [30]. In [30], the authors show that if $\mu$ is a friendly measures (see [30, §2] for definition) then any weak- limit of $\frac{-1}{\log\varepsilon}\int_{\varepsilon}^{1}\mu_{y}\,\mathrm{d}^{\times}y$ as $\varepsilon\to 0$ is a probability measure. The class of friendly measures contain the Hausdorff measures on the missing digit Cantor sets (described later). The celebrated work Simmons–Weiss [48] implies that for certain self-similar measures $\mu$, one has
$\lim_{\varepsilon\to 0}\,\frac{-1}{\log\varepsilon}\int_{\varepsilon}^{1}\mu_{y}\,\mathrm{d}^{\times}y=m_{\mathbb{X}},$ (1.2)
(in weak- sense) improving [30] in this case. Note that (1.2) is an average version of Problem 1 in its qualitative form.
In a recent breakthrough, Khalil–Luethi [27] answered Problem 1 for a class of measures that are not necessarily absolutely continuous. Below, we explain more in this direction.
1.1.1. Average Fourier decay
Our first main result is an affirmative solution of Problem 1 when $\mu$ satisfies property Item 1. We recall the definition of the $\ell^{1}$-dimension of $\mu$ (see (2.6)) as
$\dim_{\ell^{1}}\mu:=1-\inf\left\{s\geq 0\mid\lim_{X\to\infty}X^{-s}\sum_{|m|\leq X}|\widehat{\mu}(m)|=0\right\}.$
We refer to see [58, §1.4.1] and [12, §2] for more details about this quantity.
Theorem 1.
Let $\mu$ be a Borel probability measure on $\mathbb{R}$ such that
$\dim_{\ell^{1}}\mu>0.609375.$
Then there exists an $\eta>0$ and $\ell\in\mathbb{N}$ so that for any $\phi\in C_{\mathrm{b}}^{2\ell}(\mathbb{X})$ we have
$\mu_{y}(\phi)=m_{\mathbb{X}}(\phi)+O_{\mu}\left(|y|^{\eta}S_{\infty,\ell}(\phi)\right),$
as $y\to 0$.
We refer the reader to §3.3 for the definitions of $C_{\mathrm{b}}^{2\ell}(\mathbb{X})$ and $S_{\infty,\ell}(\cdot)$. The number $0.609375=\tfrac{39}{64}=\tfrac{1}{2}+\tfrac{7}{64}$ is related to the spectral gap for $\mathbb{X}$; see §3.4.3 for details.
We immediately give existence of a large class of measures that satisfy the $\ell^{1}$-dimension condition in Theorem 1. We refer to §2.2 for undefined terminologies mentioned below.
Corollary 1.
Let $\mu_{i}$ for $i=1,2$ be two Borel probability measures on $\mathbb{R}$ such that each $\mu_{i}$ is $s_{i}$-AD-regular for $s_{i}>0$. If $\frac{s_{1}+s_{2}}{2}>0.609375$ then the convolution $\mu_{1}\ast\mu_{2}$ satisfies the conclusion of Theorem 1.
Corollary 2.
Let $\mu$ be a Borel probability measure on $\mathbb{R}$ with $\dim_{\ell^{1}}\mu>0.609375$. Then for any Borel probability measure $\nu$, the convolution $\mu\ast\nu$ satisfies the conclusion of Theorem 1. Moreover, in this case the implied constant in the error term in Theorem 1 does not depend on $\nu$.
In Corollary 2, choosing $\mu$ satisfying the hypothesis of Theorem 1 and $\nu$ to be the Dirac mass $\delta_{x_{0}}$ for some $x_{0}\in\mathbb{R}$ we immediately obtain that
$\int\phi(n(x_{0}+x)g_{y})\,\mathrm{d}\mu(x)=\int_{\mathbb{X}}\phi\,\mathrm{d}m_{\mathbb{X}}+O_{\mu}(|y|^{\eta}S_{\infty,\ell}(\phi))$ (1.3)
for the same $\eta$ and $\ell$ as in Theorem 1 and uniformly in $x_{0}$.
1.1.2. Missing digit Cantor set
We define some minimal notations to state the next theorem. For $2<b\in\mathbb{N}$ and $\varnothing\neq D\subsetneq\{0,\dots,b-1\},$ let $K_{b,D}$ denote the missing digit Cantor set with base $b$ and (appearing) digits from $D$. Let $\dim_{\mathcal{H}}X$ denote the Hausdorff dimension of a set $X$. It is known that $\dim_{\mathcal{H}}K_{b,D}=\frac{\log\#D}{\log b}$; see §2.1 for details. Let $\mu_{b,D}$ denote $\frac{\log\#D}{\log b}$-dimensional Hausdorff (probability) measure restricted on $K_{b,D}$ (which we call by natural missing digit measure).
In a recent breakthrough [27], Khalil–Luethi answered Problem 1 affirmatively for any $\mu_{b,D}$ such that the Hausdorff dimension of $K_{b,D}$ is sufficiently close to $1$; in particular, $>0.9992$ for any $b$, and $>0.839$ if $b$ is prime; see [27, Theorem C, Remark 1.4] [^1] We remark that [27] considered more general rational iterated function system IFSs in higher dimensions with restrictions coming from contraction ratios and probability vectors associated to the self similar measure..
In this paper, we substantially improve upon the dimension restriction of [27].
Theorem 2.
For every $s>0.609375$, there exists explicitly computable $b(s)\in\mathbb{N}$, as in (4.3), such that for
  • any $b\geq b(s)$ and
  • any $D$ in arithmetic progression,
with
$\dim_{\mathcal{H}}K_{b,D}\geq s$
the measure $\mu_{b,D}$ satisfies the conclusion of Theorem 1.
We point out that Theorem 2 is new in this setting even in its qualitative form, i.e., without any rate of convergence. Also, we note that $b(s)$ as given in (4.3) is not optimal.
Remark 1.1.
Any $\mu_{b,D}$ with $\dim_{\mathcal{H}}K_{b,D}<0.9992$ (for general $b$) considered in Theorem 2 is not covered by [27, Theorem C]. As an easy example, let $b=450$ and $D=\{0,1,\cdots,446\}$, then $\dim_{\mathcal{H}}(K_{b,D})<0.9992$ and $\dim_{\ell^{1}}\mu_{b,D}>0.609375$ by (4.2).
Remark 1.2.
Note that the Hausdorff dimension of the middle third Cantor set (that is, $K_{3,\{0,2\}}$) is $\approx 0.6309$. Whereas in Theorem 2 we get an infinite class of natural missing digit measures whose supports have Hausdorff dimension even less than the same of middle third Cantor set. However, our result does not apply to middle third Cantor set.
Our method of proof is completely different than that of [27]; see §1.3 for details and comparison with previous methods. In particular, to prove Theorem 1, we do not need any condition (e.g. self-similarity like [27]) on the measure $\mu$ other than the restriction on $\ell^{1}$ dimension, which will be reflected in §1.1.3.
Recently, Chow–Varju–Yu in [12] find a robust way to compute $\dim_{\ell^{1}}\mu$ for self-similar measures whose attractors are missing digit Cantor sets. Thus, their method gives a large collection of non absolutely continuous measures that satisfy the hypothesis of Theorem 1.
1.1.3. Irrational IFS
For $x_{0}\in\mathbb{R}$ we consider the shifted missing digit Cantor set $K_{b,D}+x_{0}$; see (2.2). Let $\mu_{b,D,x_{0}}$ denote the $\frac{\log\#D}{\log b}$-dimensional Hausdorff measure on $K_{b,D}+x_{0}$; see (2.3).
Theorem 3.
Let $s,b,D$ be as in Theorem 2. Then for any $x_{0}\in\mathbb{R}$, the measure $\mu_{b,D,x_{0}}$ supported on $K_{b,D}+x_{0}$ satisfies the conclusion of Theorem 1.
Note that if $x_{0}$ in Theorem 3 is an irrational number then the corresponding IFS has irrational translates; see (2.2). Our next theorem, shows another instance of an IFS (in fact, uncountably many) with irrational contractions such that a self-similar measure satisfies the effective equidistribution; also see Remark 4.1.
Theorem 4.
For every sufficiently small positive $\rho$ there exists a self-similar measure with underlying IFS having (equal) contraction ratio $\rho$, satisfies the conclusion of Theorem 1.
Remark 1.3.
In [27], the authors consider IFSs that have rational contractions and translations (albeit, more general than $K_{b,D}$). We point out that when $x_{0}\notin\mathbb{Q}$ in Theorem 3 and Theorem 4 are the first instances where the Problem 1 is answered for self-similar measures associated to irrational IFSs. Moreover, in these cases the methods in [27] can not be applied as the rationality assumption of IFS is used crucially in [27].
1.1.4. Improvement assuming the Generalized Ramanujan Conjecture (GRC)
Upon the assumption of the GRC (see the discussion after (3.5)) we can replace the number $0.609375$ by $0.5$ in Theorem 1, significantly improving the $\ell^{1}$-dimension restriction. This will yield likewise improvements on e.g. Theorem 2 and Theorem 3. As an example, we describe the following corollary which follows from Theorem 1, [12, Proposition 2.4], and the proof of Theorem 3.
Corollary 3.
Assume the GRC. Let $b\geq 5$ and #$\#D=b-1$ or let $b=4$ with $D=\{0,1,2\}$ or $D=\{1,2,3\}$. Then for any $x_{0}\in\mathbb{R}$ the measure $\mu_{b,D,x_{0}}$ (see §1.1.3) satisfies the conclusions of Theorem 1.
To the best of our knowledge, albeit conditional on the GRC, Corollary 3 is the first example of missing digit Cantor sets with base $4$ where effective equidistribution, as considered above, holds.
Remark 1.4.
We end this subsection remarking that Theorem 1 on the assumption of the GRC is sharp: For every $\epsilon>0$, there exists a measure $\mu_{\epsilon}$, constructed by Kaufman in [26], that is supported on $\mathcal{W}(\psi_{\epsilon+1})$ (as defined in §1.2), such that $\dim_{\ell^{1}}\mu_{\epsilon}\geq\frac{1}{2}-\epsilon$. Thus combining [27, Theorem 9.1], the conclusion of Theorem 1 fails for $\mu_{\epsilon}$.
1.1.5. Pointwise Fourier decay
Our next main theorem answers Problem 1 for measures that satisfy Item 2.
Theorem 5.
Let $\mu$ be a Borel probability measure on $\mathbb{R}$ such that there exist a $K\in\mathbb{N}$, sequences $\{\delta_{j}\}_{j=1}^{K}\subset(0,\tfrac{1}{2}]$ and $\delta_{0}>\frac{1}{2}$, and complex numbers $\{\alpha_{j}\}_{j=1}^{K}\subset\mathbb{R}$ and $\{\beta_{j}\}_{j=1}^{K}\subset\mathbb{C}$ such that
$\widehat{\mu}(\xi)=\sum_{j=1}^{K}|\xi|^{-\delta_{j}}\beta_{j}{e(\xi\alpha_{j})}+O_{\mu}\left(|\xi|^{-\delta_{0}}\right),\quad e(z):=\exp(2\pi iz),$
as $|\xi|\to\infty$. Then there is an $\ell\in\mathbb{N}$ such that for any $0<\eta<\min\{\tfrac{1}{2},\delta_{0},\tfrac{\delta_{1}}{2}\dots,\tfrac{\delta_{K}}{2}\}$ and $\phi\in C_{\mathrm{b}}^{2\ell}(\mathbb{X})$ we have
$\mu_{y}(\phi)=m_{\mathbb{X}}(\phi)+O_{\eta,\mu}\left(|y|^{\eta}S_{\infty,\ell}(\phi)\right),$
as $y\to 0$.
Remark 1.5.
Note that in Theorem 5, the rate of Fourier decay of the measures can be arbitrarily slow, and yet we get an effective equidistribution. Moreover, in Theorem 5, in general, the measures may have $\dim_{\ell^{1}}\mu<\tfrac{1}{2}$, in fact $\dim_{\ell^{1}}\mu$ can be very close to 00. Thus Theorem 5 can not be deduced from Theorem 1.
As a proof of concept for Theorem 5, using it we answer Problem 1 for the push-forward of the Lebesgue measure by a non-constant analytic map, namely Corollary 4. However, we think that Corollary 4 can be deduced more directly (that is, not going via Theorem 5).
Let $f:\mathbb{R}\to\mathbb{R}$ be a non-constant real analytic function and $w$ be a compactly supported non-negative $L^{1}$-normalized smooth function on $\mathbb{R}$. Let $\mu^{w,f}$ be the Borel probability measure defined by $f_{\star}(w\circ\mathrm{Leb})$ where $\mathrm{Leb}$ denotes the Lebesgue measure, that is
$\mu^{w,f}(h):=\int_{\mathbb{R}}h(f(x))w(x)\,\mathrm{d}x,\quad h\in C(\mathbb{R}).$
One can check that $\mu^{w,f}$ is an absolutely continuous measure with possibly non-continuous density. One may analyse the asymptotics of the Fourier transform of $\mu^{w,f}$ via the method of stationary phase; see Proposition 5.1, which also shows that Theorem 5 is not vacuous.
Let $Z(f)$ denote the zero set of $f$. Note that as $f$ is analytic and $w$ is compactly supported the set $Z(f^{\prime})\cap\operatorname{supp}(w)$ is finite (counted with multiplicity). We define
$k_{f,w}:=\max\{\text{order of vanishing of }f^{\prime}\text{ at any }z\in Z(f^{\prime})\cap\operatorname{supp}(w)\}.$
Corollary 4.
There exists $\ell\in\mathbb{N}$ such that for any $0<\eta<\tfrac{1}{2(k_{f,w}+1)}$ and $\phi\in C_{\mathrm{b}}^{2\ell}(\mathbb{X})$ we have
$\mu^{w,f}_{y}(\phi)=m_{\mathbb{X}}(\phi)+O_{\eta}\left(|y|^{\eta}S_{\infty,\ell}(\phi)\right),$
as $y\to 0$.
One may check that the rate of equidistribution in Corollary 4 can not be improved without assuming the Riemann hypotheis (cf. the proof of [51] for $f(x)=x$). On higher rank groups the related problem is significantly difficult and, in general, open; see [46, 57, 56, 13] for recent major works.
1.2. Diophantine approximation on fractals
In this section, we focus on some applications in Diophantine approximation. Given a non-increasing monotonic positive function $\psi:\mathbb{N}\to\mathbb{R}_{+}$, we denote
$\mathcal{W}(\psi):=\{x\in[0,1]\mid|qx-p|<\psi(q)\text{ for infinitely many }q\in\mathbb{N}\text{ and }p\in\mathbb{Z}\}.$
Khintchine in 1926, [28] shows that
$\text{Leb}(\mathcal{W}(\psi))=\begin{cases}0&\text{ if }\sum\psi(q)<\infty,\\ 1&\text{ if }\sum\psi(q)=\infty.\end{cases}$ (1.4)
When $\psi_{\tau}(q):=\frac{1}{q^{\tau}}$ with $\tau>1$, the set $\bigcup_{\tau>1}\mathcal{W}(\psi_{\tau})$ is referred as the set of very well approximable numbers. Using the convergence part of Khintchine’s theorem, it follows that Lebesgue almost every point in $\mathbb{R}$ is not very well approximable.
The same observations as above for not absolutely continuous measures become highly nontrivial and are topics of research in the past two decades. For instance, in higher dimensions, studying very well approximable points inside certain manifolds is the content of Sprindžuk’s conjecture, which was famously solved by Kleinbock–Margulis in [31]. There are remarkable developments over the last two decades for general non-increasing $\psi$; see [5, 4, 2, 22, 3].
In case of measures on $\mathbb{R}$, in 1984 [35], Mahler asked “How close can irrational elements of Cantor’s (middle third) set be approximated by rational numbers not in Cantor’s set?”. Weiss in [55] shows that for the middle third Cantor set almost every point is not very well approximable with respect to the natural missing digit measure. This was significantly improved in [30], where it is shown that with respect to friendly measures almost every point is not very well approximable. On the other hand, result in [17] by Einsiedler–Fishman–Shapira, and more generally, the result in [48] by Simmons–Weiss imply that for $\psi(q)=\frac{c}{q}$ with $c>0$ one has $\mu(\mathcal{W}(\psi))=1$, when $\mu$ is a self-similar measure with an irreducible IFS satisfying the OSC in $\mathbb{R}$ (see §2.1 for definition). These results motivate the following problem which was posed by Kleinbock–Lindenstrauss–Weiss in [30]:
Problem 2.
For a friendly measure $\mu$ does the analogue of (1.4) hold?
In general, Problem 2 is wide open. In [41], the authors show under some a strong summability condition on $\psi$ together with monotonicity, with respect to absolutely friendly measure almost every point is not $\psi$ approximable.
The only known complete answer to Problem 2 is given in [27], for self-similar measures with rational irreducible IFS satisfying OSC under a certain extra condition. This extra condition in case of the natural missing digit measures (see the end of §2.1 for definition) amounts to restrict the Hausdorff dimension of the missing digit Cantor set to be very close to $1$.
With a different approach, in a groundbreaking work [58], Yu answers the convergence part of (1.4) for a class of measures $\mu$ with $\dim_{\ell^{1}}\mu>\frac{1}{2}$, namely:
Theorem.
[58, Theorem 1.5] Let $\mu$ be a Borel probability measure with $\dim_{\ell^{1}}\mu>\frac{1}{2}$. Also, let $\psi$ be a non-increasing monotonic function with $\sum\psi(q)<\infty$. Then $\mu(\mathcal{W}(\psi))=0$.
For the special function $\psi(q)=\frac{1}{q\log\log q}$, Yu in [58, Theorem 1.5] also proves divergence for self-similar measure with attractor being a missing digit Cantor set. However, his methods do not yield the divergence part in general. In this paper, we address the divergence part, extending the result of [58]. As a consequence, we extend the result of Khalil–Luethi [27, Theorem A].
Theorem 6.
Let $\mu$ be a self-similar measure whose underlying attractor is a shifted missing digit Cantor set as in §1.1.3. If
$\dim_{\ell^{1}}\mu>0.609375,$
then for any $\psi$ non-increasing monotonic function,
$\mu(\mathcal{W}(\psi))=1\quad\text{ if }\quad\sum\psi(q)=\infty.$
In particular, measures considered in Theorem 3 satisfy the above conclusion.
We remark that assuming the GRC we can improve Theorem 6 by weakening the hypothesis to $\dim_{\ell^{1}}\mu>0.5$; see Corollary 3 and discussion around it.
We note that Theorem 1 is not enough to imply Theorem 6. One needs a stronger effective equidistribution result Proposition 4.1 with any starting point, not only the identity.
We end the section remarking that when the monotonicity condition on $\psi$ is absent, an interesting question arises where approximating function is supported on a sparse set; see [34, 1, 42].
1.3. Effective equidistribution via Fourier asymptotics
We end this section with a discussion on our proofs of the effective equidistribution results.
Let $\phi\in C_{c}^{\infty}(\mathbb{X})$ with $\int_{\mathbb{X}}\phi\,\mathrm{d}m_{\mathbb{X}}=0$ and $\mu$ be a Borel probability measure on $\mathbb{R}$. Our goal is to show that $\mu_{y}(\phi)\ll|y|^{\eta}$ for some $\eta>0$. After Plancherel we write
$\mu_{y}(\phi)=\sum_{m\in\mathbb{Z}}\widehat{\phi}_{y}(m){\widehat{\mu}(m)},\quad\widehat{\phi}_{y}(m):=\int_{\mathbb{R}/\mathbb{Z}}\phi(n(x)a(y))e(-mx)\,\mathrm{d}x.$
Integrating by parts with $\left(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right)\in\mathrm{Lie}(N)$ one checks that $\widehat{\phi}_{y}(m)$ is supported essentially on $|m|\leq|y|^{-1}$.
We use Weyl’s criterion in attempt to prove the effective equidistribution. Using the spectral theory of automorphic forms on $\operatorname{SL}_{2}(\mathbb{R})$ we choose a nice basis that we plugin as $\phi$ to test the effective equidistribution. Thus for such a non-trivial automorphic form we need to essentially show that
$\sum_{|m|\leq|y|^{-1}}\widehat{\phi}_{y}(m){\widehat{\mu}(m)}\ll_{\phi}|y|^{\eta},\quad\eta>0$
with certain implicit Sobolev-type dependency on $\phi$. At this point, applying a major automorphic input, namely, theory of Whittaker function and Hecke theory we may obtain in the region $|m|\leq|y|^{-1}$ one has
$\widehat{\phi}_{y}(m)\sim\lambda(|m|)|y|^{1/2}$
where $\lambda(m)$ are the Hecke eigenvalues. The GRC predicts that $\lambda(m)$ are essentially bounded, where as unconditionally, by the spectral gap [29] one has $\lambda(m)\ll m^{\vartheta}$ for any $\vartheta>\frac{7}{64}$. Thus to show effective equidistrbution we need to show a better than square-root cancellation, namely,
$\sum_{m\leq X}\lambda(m)\widehat{\mu}(m)\ll X^{\frac{1}{2}-\eta},\quad\eta>0,X\to\infty.$
Now we list down what can say about the above sum using different techniques:
  • Trivial bound: Note that trivially $|\widehat{\mu}(m)|\leq 1$. Thus assuming the GRC the above sum is bounded by $X$ (unconditionally, it is even worse).
  • Cauchy–Schwarz and average GRC: After Cauchy–Schwarz and applying (3.8) and (2.6) one can bound the sum by
    $\sqrt{\sum_{m\leq X}|\lambda(m)|^{2}}\sqrt{\sum_{m\leq X}|\widehat{\mu}(m)|^{2}}\ll X^{\frac{1}{2}+\frac{1-\dim_{\ell^{2}}\mu}{2}}.$
    To win, one needs $\dim_{\ell^{2}}\mu>1$ which is impossible!
Remark 1.6.
We briefly remark for a missing digit Cantor set the above approach is somewhat close to that of Khalil–Luethi [27], where they start with a self-similar $\mu$ (albeit, they have not used a Fourier dualization to start with). They used self-similarity quite cleverly and crucially to start with $\mu_{y}(\mathcal{P}^{n}\phi)$ where $\mathcal{P}$ is the Markov operator (see [27, Definition 2.1]) underlying the self-similar structure of $\mu$. Heuristically, this replaces $\lambda(m)$ in the sum in consideration by the “Hecke eigenvalues”, say $\nu_{n}(m)$ of $\mathcal{P}^{n}\phi$. Their winning point can be summaraized as that such $\nu_{n}$ has a “better spectral gap on average”.
Now we describe the winning strategy that we take in this paper.
  • Hölder estimate with $(1,\infty)$-pair: We apply Hölder’s estimate on the sum in consideration. Applying the bound of Hecke eigenvalues we see the sum is
    $\ll X^{\vartheta+1-\dim_{\ell^{1}}\mu}=X^{\frac{1}{2}-\left(\dim_{\ell^{1}}\mu-\frac{1}{2}-\vartheta\right)},$
    which is the main idea behind our Theorem 1. We remark that this argument follows without any self-similarity assumption on $\mu$, unlike [27].
  • Fourier decay of $\mu$: Assume that $\widehat{\mu}(m)\ll m^{-\delta}$ for some $\delta>0$. Then using average GRC we can see that equidistribution holds if $\delta>\frac{1}{2}$; (cf. Lemma 5.1).
  • Cancellation of additive twist: A deep automorphic result of Jutila [25] gives us
    $\sum_{m\leq X}\lambda(m)e(mx)\ll X^{\frac{1}{2}},\quad\text{uniformly in }x,$
    which immediately implies that the sum in consideration is $O(X^{\frac{1}{2}})$. This method fails but almost at the border. An important point is that to reach at the border we have not assumed anything on $\mu$.
    The above gives us a hint that if $\widehat{\mu}(m)$ asymptotoically behaves as $e(m\alpha)m^{-\delta}$ for any $\delta>0$ we have a chance of winning. This and the previous point are the main ideas behind our Theorem 5.
2. Measure Theoretic Preliminaries
2.1. Self-similar measures
We record certain basic definitions about self-similar sets and measures in $\mathbb{R}$. We refer to [38] for details.
Let $l\in\mathbb{N}$ and =$\mathcal{F}=\{f_{i}\}_{1\leq i\leq l}$ be a finite collection contracting similarities, i.e.,
$f_{i}:\mathbb{R}\to\mathbb{R},\quad x\mapsto\rho_{i}x+\theta_{i},\quad 0<\rho_{i}<1,\quad\theta_{i}\in\mathbb{R}.$
Such an $\mathcal{F}$ is called an Iterated Function System (IFS). Using Banach fixed point theorem (see [23]), we see that there exists a unique compact set $K\subset\mathbb{R}$, known as attractor, such that
$\bigcup_{1\leq i\leq l}f_{i}(K)=K.$
We say $\mathcal{F}$ satisfies the Open Set Condition (OSC) if there exists an open set $U\subset\mathbb{R}$ such that
$f_{i}(U)\subseteq U,\quad f_{i}(U)\cap f_{j}(U)=\varnothing,\quad 1\leq i\neq j\leq l.$
In this case, the Hausdorff dimension of $K$, denoted by $\dim_{\mathcal{H}}K$, is the unique solution $s$ of the equation
$\sum_{i=1}^{n}\rho_{j}^{s}=1.$
Finally, we call $\mathcal{F}$ to be irreducible if there does not exist an $\alpha\in\mathbb{R}$ such that $f_{i}(\alpha)=\alpha$ for $1\leq i\leq l$.
Let $\lambda=\{\lambda_{i}\}_{1\leq i\leq l}$ be a probability vector, i.e.,
$\sum_{i=1}^{l}\lambda_{i}=1,\quad\lambda_{i}>0.$
Given an IFS $\mathcal{F}$ and probability vector $\lambda$ there exists an unique Borel probability measure $\mu$ such that
$\sum_{i=1}^{l}\lambda_{i}{f_{i}}_{\star}\mu=\mu,$ (2.1)
where by $f_{\star}\mu$ we denote the push-forward of $\mu$ by a function $f$. Such a measure is called self-similar measure. If $\mathcal{F}$ satisfies OSC then $\mu$ has null overlaps, that is,
$\mu(f_{i}(K)\cap f_{j}(K))=0,\quad 1\leq i\neq j\leq l,$
see [20, Lemma 3.3]. If $\mathcal{F}$ has OSC, then for $\lambda_{j}=\rho_{j}^{s}$ the corresponding self-similar measure $\mu$ is the $s$-dimensional Hausdorff measure restricted to the attractor $K$ which is $s$-AD regular; see [37, Theorem 4.14].
Let $b\geq 3$ and $\varnothing\neq D\subset\{0,\cdots,b-1\}$ with #$\#D=l\geq 2$. Let $\mathcal{F}_{b,D}$ denote the IFS $\{f_{i}(x):=\frac{x+i}{b}\mid i\in D\}$. The IFS $\mathcal{F}_{b,D}$ satisfies OSC and the attractor of $\mathcal{F}_{b,D}$ is $K_{b,D}$. Given any probability vector, there exists an unique self-similar measure whose attractor is $K_{b,D}$. We refer to them as self-similar measures for which the attractor is a missing digit Cantor set. In particular, when the probability vector is uniform probability vector, we call the self-similar measure to be natural missing digit measure $\mu_{b,D}.$ Note $\mu_{b,D}$ is also the $s=\frac{\log l}{\log b}$-dimensional Hausdorff measure restricted on $K_{b,D}.$
For $x_{0}\in\mathbb{R}$ it can be easily checked that $K_{b,D}+x_{0}$ is the attractor of the IFS
$\mathcal{F}_{b,D,x_{0}}:=\left\{f_{i}(x):=\frac{x+i}{b}+x_{0}(1-\tfrac{1}{b})\mid i\in D\right\}$ (2.2)
and the convolution
$\mu_{b,D,x_{0}}:=\mu_{d,D}\ast\delta_{x_{0}}$ (2.3)
is the self-similar measure on $K_{b,D}+x_{0}$ associated with uniform probability vector. It is easy to check that $\mathcal{F}_{b,D,x_{0}}$ satisfies OSC.
2.2. Fourier theory
Let $\mu$ be a Borel probability measure on $\mathbb{R}$. Given two Borel probability measures $\mu_{1}$ and $\mu_{2}$ we denote their convolution by
$\mu_{1}\ast\mu_{2}(f)=\int f(x+y)\,\mathrm{d}\mu_{1}(x)\,\mathrm{d}\mu_{2}(y),\quad f\in C(\mathbb{R}).$
We define the Fourier transform of $\mu$ by
$\widehat{\mu}(\xi):=\int e(\xi x)\,\mathrm{d}\mu(x);\quad e(z):=\exp(2\pi iz),z\in\mathbb{C}.$ (2.4)
If $\mu$ is compactly supported then $\widehat{\mu}$ is a bounded Lipschitz continuous function. Note that $\widehat{\mu_{1}\ast\mu_{2}}=\widehat{\mu_{1}}\widehat{\mu_{2}}$. It follows from the classical Riemann–Lebesgue lemma that any absolute continuous measure (w.r.t the Lebesgue measure) $\widehat{\mu}(\xi)\to 0$ as $|\xi|\to\infty$. However, it is often extremely difficult to quantify the decay rate of $\widehat{\mu}$ in general.
On the other hand, when $\mu$ is not absolutely continuous, $\widehat{\mu}$ may not have any decay. For instance, if $\mu$ is the Hausdorff measure supported on the middle-third Cantor set then $\widehat{\mu}(\xi)\nrightarrow 0$, i.e., $\widehat{\mu}$ has no decay. If $\mu$ is a self-similar measure (see §2.1) with underlying IFS having $\rho_{i}=\rho$ for $1\leq i\leq l$ then it is easy to compute the Fourier transform of $\mu$, namely
$\widehat{\mu}(\xi)=\prod_{j=1}^{\infty}g(\rho^{j}\xi),\quad g(\xi):=\sum_{i=1}^{l}\lambda_{i}e(b_{i}\xi);$ (2.5)
see [50, Lemma 4.1]. We refer readers to [54, 43, 47] for surveys on behaviour of Fourier transform in case of various self-similar measures.
In this paper, we mainly consider the Fourier decay in an $\ell^{p}$-average sense. Following [58, Definition 1.2], we define,
$\dim_{\ell^{p}}\mu:=1-\inf\left\{\ell\mid\lim_{X\to\infty}X^{-\ell}\sum_{|m|\leq X}|\widehat{\mu}(m)|^{p}=0\right\}.$ (2.6)
The set on the right-hand side above is non-empty as trivially any $\ell>1$ belongs to this set. Moreover, this definition is equivalent to those in [58, Definition 1.2]. Indeed, this follows from the fact for any sequence $\{A(X)\}_{X\in\mathbb{N}}$ of positive reals
$A(X)=o_{\epsilon}(X^{\epsilon})\,\forall\epsilon>0\iff A(X)=O_{\epsilon}(X^{\epsilon})\,\forall\epsilon>0.$
Using Cauchy–Schwarz inequality and definitions, as noted in [58, Lemma 1.4], we have
$\frac{1}{2}\dim_{l^{2}}\mu\leq\dim_{\ell^{1}}\mu\leq\dim_{l^{2}}\mu.$
For $s>0$ we call a Borel measure $\mu$ to be $s$-AD regular if there is a $C>1$ such that for any $x\in\operatorname{supp}(\mu)$ and sufficiently small $r>0$ one has
$C^{-1}r^{s}\leq\mu((x-r,x+r))\leq Cr^{s}.$
By [38, §3.8], if $\mu$ is $s$-AD regular then,
$\dim_{\ell^{2}}\mu=s.$ (2.7)
We define a variant of the above $\ell^{p}$ dimension, following [59, Definition 2.2]. For $1\leq p<\infty$ we define
$\dim^{\star}_{\ell^{p}}\mu:=1-\inf\left\{\ell\mid\lim_{X\to\infty}X^{-\ell}\sup_{0\leq\theta\leq 1}\sum_{|m|\leq X}|\widehat{\mu}(m+\theta)|^{p}=0\right\}.$ (2.8)
It is immediate to check that
$\dim^{\star}_{\ell^{p}}\mu\leq\dim_{\ell^{p}}\mu.$ (2.9)
However, for $p=1$ a self-similar measure whose attractor is a missing digit Cantor set (see 2.1), the above two dimensions will coincide; see Lemma 4.4.
3. Automorphic Theory
3.1. Convention
The letter $\epsilon$ will denote placeholders for a small positive number. In particular, the exact value of $\epsilon$ may change from line to line. We will make the same convention with the order of the Sobolev norm (denoted by $d$ or $\ell$) which is thought as an implicit positive integer. Moreover, we will write $\ll x^{O_{P}(1)}$ to denote $\ll x^{d}$ for some unspecified positive $d$ depending on $P$ as $x\to\infty$.
3.2. Groups and measures
Let $G:=\operatorname{PGL}_{2}(\mathbb{R})$, $\Gamma:=\operatorname{PGL}_{2}(\mathbb{Z})$ and $[G]:=\Gamma\backslash G$. Throughout the paper we will identify $\mathbb{X}$ with $[G]$ by means of, e.g., [27, Lemma 3.1]. We denote
$n(x):=\begin{pmatrix}1&x\\ &1\end{pmatrix},\quad a(y):=\begin{pmatrix}y&\\ &1\end{pmatrix}.$
Via the above identification we rewrite $\mu_{y}$, as in (1.1), for a Borel probability measure $\mu$ as
$\mu_{y}(\phi)=\int\phi(n(x)a(y))\,\mathrm{d}\mu(x),\quad\phi\in C(\mathbb{X}).$
Let $K:=\mathrm{SO}_{2}(\mathbb{R})$, which is a compact subgroup of $G$. We will abbreviate $\exp(2\pi iz)$ by $e(z)$ for any $z\in\mathbb{C}$.
We fix measures $\,\mathrm{d}x$ on $N:=\{n(x)\mid x\in\mathbb{R}\}$ and $\,\mathrm{d}^{\times}y:=\frac{\,\mathrm{d}y}{|y|}$ on $A:=\{a(y)\mid y\in\mathbb{R}^{\times}\}$. Correspondingly, we can write a Haar measure on $G$ in the Iwasawa coordinates $g=n(x)a(y)k$ as $\,\mathrm{d}g=\,\mathrm{d}x\frac{\,\mathrm{d}^{\times}y}{|y|}\,\mathrm{d}k$ where $\,\mathrm{d}k$ is the probability Haar measure on $K$. We use $m_{[G]}$ (which is the same is $m_{\mathbb{X}}$) to denote the invariant probability measures on on $[G]$.
3.3. Sobolev norm
We refer to [10, Chapter 2] for details. Let $\{X,Y,Z\}$ be an orthonormal basis of $\mathrm{Lie}(G)$ where the orthonormality is with respect to the Killing form on $\mathrm{Lie}(G)$ and $Z$ is a non-zero element in $\mathrm{Lie}(K)$. We define a Laplacian of $G$ by the formula
$\mathcal{D}:=\mathrm{Id}-X^{2}-Y^{2}-Z^{2}=\mathrm{Id}-C_{G}+2C_{K}$ (3.1)
where $C_{G}:=X^{2}+Y^{2}-Z^{2}$ is the Casimir operator on $G$ and $C_{K}:=-Z^{2}$ is the Casimir operator on $K$
Let $\pi$ be a unitary representation of $G$. It is known that (see, e.g., [40, proof of Theorem 3]) $\mathcal{D}|_{\pi}$ is a densely-defined, self-adjoint, invertible, positive definite, second-order elliptic differential operator. For $d\in\mathbb{Z}_{\geq 0}$ we define the $d$’th Sobolev norm on $\pi$ by
$S_{2,d}(v):=\|\mathcal{D}^{d}v\|_{\pi},\quad\text{smooth }v\in\pi$
and extend the definition to $\pi$ by density of the smooth vectors in $\pi$.
The Casimir operator $C_{G}$ lies in the center of the universal enveloping algebra of $G$. Thus if $\pi$ is also irreducible then $C_{G}$ acts on $\pi$ by a scalar $\nu_{\pi}$. On the other hand, if $v\in\pi$ is a $K$-type vector (i.e. $K$ acts on $v$ by a character) of weight $k\in\mathbb{Z}$ then $C_{K}$ also acts on $v$ by a scalar $\nu_{k}$. Consequently, it follows from (3.1) that $\mathcal{D}$ acts on a weight-$k$ vector $v\in\pi$ by a scalar. Thus a $\mathcal{D}$-eigenbasis of $\pi$ is the same as a $K$-isotypic basis of $\pi$.
Lemma 3.1.
For every $\ell>0$ there is a $\ell^{\prime}>0$ such that we have
$\mathrm{trace}\left(\mathcal{D}^{-\ell^{\prime}}|_{\pi}\right)\ll(1+|\nu_{\pi}|)^{-\ell}$
for any unitary irreducible $\pi$.
Proof.
Let $\{v_{k}:k\in\mathbb{Z}\}$ be an orthonormal $K$-type basis of $\pi$ where $v_{k}$ is a weight $k$-vector. Thus, from the above discussion it follows that
$\mathrm{trace}\left(\mathcal{D}^{-\ell^{\prime}}|_{\pi}\right)=\sum_{k\in\mathbb{Z}}(1-\nu_{\pi}+2\nu_{k})^{-\ell^{\prime}}.$
From [10, Proposition 2.5.4], we have that $\nu_{k}\asymp k^{2}$. Thus the above sum is bounded by
$\ll\sum_{k\in\mathbb{Z}}(1+|\nu_{\pi}|)^{-\frac{\ell^{\prime}}{2}}(1+|k|)^{-\ell^{\prime}}.$
Making $\ell^{\prime}$ sufficiently large we conclude. ∎
Finally, for $0\leq\ell\leq\infty$, we define
$C_{\mathrm{b}}^{2\ell}([G]):=\{\phi\in C^{2\ell}([G])\mid S_{\infty,\ell}(\phi):=\|\mathcal{D}^{\ell}\phi\|_{L^{\infty}([G])}<\infty\}.$
Note that for every $\ell>0$ and for every $\phi\in C_{\mathrm{b}}^{2\ell}([G])$ we have
$S_{\infty,\ell}(\phi)\ll\mathcal{S}_{\infty,\ell}(\phi)$
where $\mathcal{S}_{\infty,\ell}$ is the Sobolev norm defined in [27, eq.(3.6)].
3.4. Automorphic forms and representations
We refer to [14] for details. In this section, $\pi$ will denote a standard non-trivial unitary automorphic representation for $\Gamma$, that is, $\pi$ is infinite dimensional and appears in the spectral decomposition of $L^{2}([G])$; see Lemma 3.5. Then $\pi$ is either cuspidal or a unitary Eisenstein series. By $\varphi\in\pi$ (cuspidal or an Eisenstein series) we will denote an automorphic form on $[G]$.
3.4.1. Unitary structure
If $\pi$ is cuspidal then we fix a $G$-invariant inner product on $\pi$ by
$\|\varphi\|^{2}_{\pi}:=\|\varphi\|^{2}_{[G]}=\int_{[G]}|\varphi(g)|^{2}\,\mathrm{d}g.$
If $\pi$ is unitary Eisenstein then any $\varphi\in\pi$ is of the form
$\varphi=\operatorname{Eis}(f),\quad f\in\mathcal{I}(s):=\mathrm{Ind}_{NA}^{G}|\cdot|^{s},\quad s\in i\mathbb{R};$
where for $\Re(s)>\tfrac{1}{2}$ we define
$\operatorname{Eis}(f):=\sum_{\gamma\in N\cap\Gamma\backslash\Gamma}f(\gamma\cdot).$
The sum converges absolutely for $\Re(s)>\tfrac{1}{2}$ and can be meromorphically continued to all $s\in\mathbb{C}$. Note that $f$ satisfies
$f(n(x)a(y)g)=|y|^{\frac{1}{2}+s}f(g),\quad x\in\mathbb{R},y\in\mathbb{R}^{\times},g\in G.$ (3.2)
Thus any such $f$ is determined by $f|_{K}$ by Iwasawa decomposition. If $s\in i\mathbb{R}$ then $\mathcal{I}(s)$ is unitary. In this case, we fix a $G$-invariant inner product on $\mathcal{I}(s)$ by
$\|f\|^{2}_{\mathcal{I}(s)}:=\int_{K}|f(k)|^{2}\,\mathrm{d}k$
and a $G$-invariant inner product on $\pi$ by
$\|\operatorname{Eis}(f)\|_{\pi}=\|f\|_{\mathcal{I}(s)}.$
Through out the paper we only work with $f\in\mathcal{I}(s)$ such that $f|_{K}$ is independent of $s$. In this paper, if not mentioned otherwise, we assume that $f|_{K}$ is $s$-independent and a $K$-type vector. In this case, $\operatorname{Eis}(f)$ is holomorphic in the region $0\leq\Re(s)<\tfrac{1}{2}$.
3.4.2. Constant term
We define the constant term of $\varphi$ by
$\varphi_{0}(g):=\int_{\mathbb{R}/\mathbb{Z}}\varphi(n(x)g)\,\mathrm{d}x.$
If $\varphi$ is cuspidal then $\varphi_{0}=0$. If $\varphi=\operatorname{Eis}(f)$ then (see [24, eq.(6)])
$\varphi_{0}(g)=f(g)+\frac{\zeta(2s)}{\zeta(1+2s)}M(s)f(g),$ (3.3)
where $\zeta$ is the Riemann zeta function and $M(s)$ is the standard intertwiner mapping $\mathcal{I}(s)\to\mathcal{I}(-s)$, defined by
$M(s)f(g):=\int_{\mathbb{R}}f(wn(x)g)\,\mathrm{d}x,\quad w:=\begin{pmatrix}&1\\ 1&\end{pmatrix}.$
The above integral converges absolutely for $\Re(s)>0$ and has meromorphic continuation to all of $\mathbb{C}$.
Lemma 3.2.
Let $\varphi=\operatorname{Eis}(f)$ for $f\in\mathcal{I}(s)$. Then for all $x\in\mathbb{R}$, $y\in\mathbb{R}^{\times}$, and $k\in K$ we have
$\varphi_{0}(n(x)a(y)k)=\varphi_{0}(a(y)k)=|y|^{\frac{1}{2}}\left(|y|^{s}f(k)+\frac{\zeta(2s)}{\zeta(1+2s)}|y|^{-s}M(s)f(k)\right).$
In particular, there is a $d>0$ so that for $\Re(s)=0$
$\varphi_{0}(n(x)a(y)k)\ll|y|^{\frac{1}{2}}S_{2,d}(\varphi)$
uniformly in $x$ and $k$.
Proof.
The formula of $\varphi_{0}$ and left $N$-invariance of it follow immediately from (3.3), and the facts that $f\in\mathcal{I}(s)$ and consequently, $M(s)f\in\mathcal{I}(-s)$. To see the next estimate, we first use the functional equation of the Riemann zeta function:
$\frac{\zeta(2s)}{\zeta(1+2s)}=\frac{\Gamma_{\mathbb{R}}(1-2s)}{\Gamma_{\mathbb{R}}(2s)}\frac{\zeta(1-2s)}{\zeta(1+2s)},\quad\Gamma_{\mathbb{R}}(s):=\pi^{-s/2}\Gamma(s/2).$
Note that
$\left|\frac{\zeta(1-2s)}{\zeta(1+2s)}\right|=1,\quad\Re(s)=0.$
Then the claim follows from the fact that
$f(k),\quad\frac{\Gamma_{\mathbb{R}}(1-2s)}{\Gamma_{\mathbb{R}}(2s)}M(s)f(k)\ll S_{2,d}(\varphi);$
see [24, eq.(7), §2.2] and the discussion around it. ∎
3.4.3. Fourier coefficients
For any cuspidal or unitary Eisenstein series $\pi\ni\varphi$, one has the Fourier expansion
$\varphi(g)-\varphi_{0}(g)=\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{\sqrt{|m|}}W_{\varphi}(a(m)g).$ (3.4)
Let $\pi$ be cuspidal and $\varphi\in\pi$. In this paper, we always use the normalized $\varphi$ with $\lambda_{\pi}(1)=1$. In this case, $\lambda_{\pi}$ are the Hecke eigenvalues attached to $\pi$. The Whittaker function $W_{\varphi}$ in this case is given by
$W_{\varphi}(g):=\int_{\mathbb{R}/\mathbb{Z}}\varphi(n(x)g)e(-x)\,\mathrm{d}x.$
The expression (3.4) follows from Shalika’s multiplicity one theorem; see [14, §4.1 and §6.2]. We have the following bound
$\lambda_{\pi}(|m|)\ll|m|^{\vartheta},\quad\vartheta>\frac{7}{64},$ (3.5)
uniformly in $\pi$; see [29, Proposition 2]. The Generalized Ramanujan Conjecture (GRC) predicts that the above estimate of $\lambda_{\pi}$ can be improved to $|m|^{\epsilon}$ for any $\epsilon>0$. However, this is only known when $\pi$ corresponds to a modular form (that is, a discrete series representation).
If $\pi$ is an Eisenstein series with parameter $s\in\mathbb{C}$ and $\varphi=\operatorname{Eis}(f)$ then the Whittaker function $W_{\varphi}$ is given by
$W_{\varphi}(g):=W_{f}(g):=\int_{\mathbb{R}}f(wn(x)g)e(-x)\,\mathrm{d}x$
which converges absolutely for $\Re(s)>0$ and can be analytically continued to all of $\mathbb{C}$. In particular, when $f$ varies over a family such that $f|_{K}$ is $s$-independent (so called flat family) then $W_{f}$ is entire in $s$. In this case, we have
$\lambda_{\pi}(|m|)=\frac{|m|^{-s}\tau_{2s}(|m|)}{\zeta(1+2s)},\quad\tau_{z}(m):=\sum_{d\mid m}d^{z};$ (3.6)
see [24, §2.1] for details of the proof. Consequently, for $\Re(s)=0$ it follows from the lower bound of $\zeta$ on $\Re(s)=1$ that (see [53, pp.50-51] and [10, Proposition 2.5.4])
$\lambda_{\pi}(|m|)\ll_{\epsilon}\left(|m|(1+|s|)\right)^{\epsilon}\asymp\left(|m|(1+|\nu_{\pi}|)\right)^{\epsilon}.$ (3.7)
Finally, the GRC is known on average for any cuspidal or unitary Eisenstein series. That is,
$\sum_{m\leq X}|\lambda_{\pi}(m)|^{2}\ll_{\epsilon}X^{1+\epsilon}(1+|\nu_{\pi}|)^{\epsilon};$ (3.8)
see [16, Proposition 19.6].
3.4.4. Functional equation of the Eisenstein series
We record the functional equation of the Eisenstein series. It follows from (see [8, eq.(3.10)], cf. [10, Proposition 4.5.9]) that
$W_{f}=\frac{\Gamma_{\mathbb{R}}(1-2s)}{\Gamma_{\mathbb{R}}(2s)}W_{M(s)f}=W_{M^{\ast}(s)f},\quad M^{\ast}(s):=\frac{\Gamma_{\mathbb{R}}(1-2s)}{\Gamma_{\mathbb{R}}(2s)}\dot{M}(s).$ (3.9)
Moreover, by Schur’s lemma and the above it follows that
$M^{\ast}(-s)\circ M^{\ast}(s)=\mathrm{Id}.$
Thus from the Fourier expansion of Eisenstein series, namely, (3.4), (3.3), and (3.6), and the functional equation of the Riemann zeta function it follows that
$\zeta(1+2s)f(g)+\zeta(2s)M(s)f(g)+\sum_{m\neq 0}\frac{|m|^{-s}\tau_{2s}(|m|)}{\sqrt{|m|}}W_{f}(a(m)g)\\ =\zeta(1-2s)M^{\ast}(s)f(g)+\zeta(-2s)M(-s)\circ M^{\ast}(s)f(g)+\sum_{m\neq 0}\frac{|m|^{s}\tau_{-2s}(|m|)}{\sqrt{|m|}}W_{M^{\ast}(s)f}(a(m)g)$
Hence, we obtain
$\zeta(1+2s)\operatorname{Eis}(f)=\zeta(1-2s)\operatorname{Eis}(M^{\ast}(s)f)=\zeta(2s)\operatorname{Eis}(M(s)f)$ (3.10)
for any $s\in\mathbb{C}$.
3.5. Whittaker functions
In this section, $\pi$ will denote an abstract unitary irreducible representation of $G$. We call $\pi$ to be generic if $\pi$ has a unique $G$-invariant embedding into
[…]
$W(n(x)g)=e(x)W(g),\quad\forall x\in\mathbb{R},g\in G.$ (3.11)
The functions $W$ are known as Whittaker functions and the image of $\pi$ inside $\mathrm{Ind}_{N}^{G}\,e(\cdot)$ under the above embedding is known as the Whittaker model of $\pi$. In this paper, we always identify a generic representation and its Whittaker model.
3.5.1. Mellin Theory
From the theory of local $\operatorname{GL}(2)\times\operatorname{GL}(1)$ Hecke zeta integral (see, e.g., [14, Chapter 8]) we know that for $\varepsilon\in\{0,1\}$ and $s\in\mathbb{C}$ the zeta integral
$Z(s,\varepsilon,W):=\int_{\mathbb{R}^{\times}}W(a(y))|y|^{s}\operatorname{sgn}(y)^{\varepsilon}\,\mathrm{d}^{\times}y$ (3.12)
converges absolutely for $\Re(s)>-\frac{1}{2}$ and hence defines a holomorphic function in this region. The same also follows from the growth of $W(a(y))$, namely,
$(y\partial_{y})^{j}W(a(y)k)\ll_{j,A}|y|^{-A}S_{2,d}(W),\quad y\in\mathbb{R}^{\times},k\in K;$ (3.13)
where $A$ can take any value in $(-\frac{1}{2},\infty)$. Here $d$ only depends on $j$ and $A$. The above follows from [39, Proposition 3.2.3] temperedness of $\pi$ (at the archimedean place) due to Selberg. Finally, we record the functional equation of the local $\operatorname{GL}(2)\times\operatorname{GL}(1)$ zeta integral (see [14, eq.(8.1), Proposition 8.1])
$Z(s,\varepsilon,W)=Z(-s,\varepsilon,W(\cdot w))\gamma(\tfrac{1}{2}-s,\pi)$ (3.14)
where $\gamma$ is a certain complex meromorphic function (implicitly depending on $\varepsilon$) satisfying
$\gamma(\tfrac{1}{2}-s,\pi)\ll_{\Re(s)}(1+|s|)^{2\Re(s)}\nu_{\pi}^{O(1)}$ (3.15)
as long as $s$ is a fixed distance away from any pole of the $\gamma$-function; see [14, Proposition 9.5–9.6].
3.5.2. Unitary structure
We fix a $G$-invariant unitary structure on the Whittaker model of $\pi$ by
$\|W\|^{2}_{\pi}:=\int_{\mathbb{R}^{\times}}|W(a(y)|^{2}\,\mathrm{d}^{\times}y.$
It is known that cuspidal representations and unitary Eisenstein series $\pi$ of $G$ are generic. For any such $\varphi\in\pi$ with Whittaker function $W_{\varphi}$ it follows from Schur’s lemma that (see [14, §5.2-5.3])
$\|\varphi\|^{2}_{\pi}=C_{\pi}\|W_{\varphi}\|^{2}_{\pi}.$
If $\pi$ is cuspidal then from (3.8) and [21, Theorem 0.2] it follows that
$C_{\pi}^{-1}\ll_{\epsilon}(1+|\nu_{\pi}|)^{\epsilon}$
for any $\epsilon>0$.
If $\pi$ is a unitary Eisenstein series with parameter $s\in i\mathbb{R}$ then $C_{\pi}=1$ which follows from the fact that
$\|W_{f}\|_{\mathcal{I}(s)}=\|f\|_{\mathcal{I}(s)};$
see, e.g., [8, eq.(3.11)]. The above combining with (3.9) implies that
$\|f\|_{\mathcal{I}(s)}=\|M^{\ast}(s)f\|_{\mathcal{I}(-s)}.$ (3.16)
Finally, if $\varphi$, cuspidal or unitary Eisenstein, is a $\mathcal{D}$-eigenfunction if and only if $W_{\varphi}$ is the same with the same eigenvalue. Thus the above estimates of $C_{\pi}$ imply that
$S_{2,d}(W_{\varphi})\ll_{\epsilon}(1+|\nu_{\pi}|)^{\epsilon}S_{2,d}(\varphi)\ll S_{2,d}(\varphi)$
for any $\mathcal{D}$-eigenfunction $\varphi$. From now on, if not mentioned otherwise, we assume that $\varphi$ is a $\mathcal{D}$-eigenfunction and use the above without anymore explanation.
Lemma 3.3 (Truncation).
Fix $\sigma>1$. For every $N>0$ sufficiently large there is a $d>0$ so that for every automorphic form $\varphi$ we have
$(\varphi-\varphi_{0})(n(x)a(y))=\sum_{0\neq|m|\leq|y|^{-\sigma}}\frac{\lambda_{\pi}(|m|)}{\sqrt{|m|}}W_{\varphi}(a(my))e(mx)+O\left(|y|^{N}S_{2,d}(\varphi)\right)$
as $y\to 0$.
Proof.
We start with the Fourier–Whittaker expansion of $\varphi$ is in (3.4). Using unipotent equivariance of the Whittaker function as in (3.11) we obtain
$(\varphi-\varphi_{0})(n(x)a(y))=\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{\sqrt{|m|}}W_{\varphi}(a(my))e(mx).$
Thus it is enough to show that the contribution of the terms for $|m|>|y|^{-\sigma}$ in the above sum is $O\left(|y|^{N}S_{2,d}(\varphi)\right)$ for all large $N$.
We recall from (3.13) the bound of the Whittaker function
$W_{\varphi}(a(my))\ll_{A}|my|^{-A}S_{2,d}(\varphi).$
From (3.5) and (3.7) we obtain the bound of Hecke eigenvalues
$\frac{\lambda_{\pi}(|m|)}{\sqrt{|m|}}\ll(1+|\nu_{\pi}|)^{\epsilon},\quad m\neq 0.$
Thus using triangle inequality and partial summation we obtain
$\sum_{|m|\geq|y|^{-\sigma}}\frac{\lambda_{\pi}(|m|)}{\sqrt{|m|}}W_{\varphi}(a(my))e(mx)\ll S_{2,d}(\varphi)\sum_{|m|\geq|y|^{-\sigma}}|my|^{-A}\ll|y|^{A(\sigma-1)-\sigma}S_{2,d}(\varphi).$
Making $A$ sufficiently large we conclude. ∎
Lemma 3.4.
There is a $d>0$ such that for any unitary automorphic form $\varphi$ we have
$\|\varphi\|_{1}\ll S_{2,d}(\varphi)$
and
$\varphi(n(x)a(y)k)\ll_{\epsilon}\max\left(\sqrt{|y|}\,,\,|y|^{-1-\epsilon}\right)S_{2,d}(\varphi)$
uniformly in $x\in\mathbb{R}$.
The second assertion is essentially the Sobolev embedding as stated in [39, S2a]. However, we give a proof for completion.
Proof.
Applying the Fourier–Whittaker expansion (3.4) and unipotent equivariance of Whittaker function (3.11) for any $g=n(x)a(y)k$ we write
$(\varphi-\varphi_{0})(g)=\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{\sqrt{|m|}}W_{\varphi}(a(my)k)e(mx).$
Using the bounds (3.5), (3.7), (3.13) and working as in the proof of Lemma 3.3 with $A=1+\epsilon$ we bound the above by
$S_{2,d}(\varphi)\sum_{m\neq 0}|my|^{-1-\epsilon}\ll|y|^{-1-\epsilon}S_{2,d}(\varphi)$
for some fixed $d>0$. Finally, applying Lemma 3.2 we obtain
$\varphi(g)\ll\max\left(\sqrt{|y|}\,,\,|y|^{-1-\epsilon}\right)S_{2,d}(\varphi),\quad g=n(x)a(y)k.$
This proves the second assertion.
To prove the first assertion we fix the Siegel domain
$S:=\{g=n(x)a(y)k\mid|x|\leq\tfrac{1}{2},\,y\geq\tfrac{\sqrt{3}}{2},\,k\in K\}\supseteq[G].$
Then we obtain
$\|\varphi\|_{1}\leq\int_{S}|\varphi|\ll S_{2,d}(\varphi)\int_{|x|\leq\frac{1}{2}}\,\mathrm{d}x\int_{|y|\gg 1}\sqrt{|y|}\frac{\,\mathrm{d}^{\times}y}{|y|}\int_{K}\,\mathrm{d}k\ll S_{2,d}(\varphi)$
as claimed. ∎
Remark 3.1.
For $0\leq\Re(s)<\tfrac{1}{2}$ and $f\in\mathcal{I}(s)$ it follows from the proof of Lemma 3.4 and (3.3) that $\operatorname{Eis}(f)$ is $L^{1}$-integrable on $[G]$.
3.6. Spectral decomposition
For a unitary representation $\pi$ (similarly, $\mathcal{I}(s)$) of $G$ by $\mathcal{B}(\pi)$ we denote an orthogonal basis of $\pi$ consisting of arithmetically normalized $K$-type vectors. For the rest of the paper, by $\mathcal{B}(\pi)$ we mean a basis of $\pi$ with the properties as above, unless otherwise mentioned.
Lemma 3.5.
There exist a sufficiently large $\ell>0$ such that for any $\phi\in C_{\mathrm{b}}^{2\ell}([G])$ and any $g\in G$ we have
$\phi(g)=\int_{[G]}\phi\,\mathrm{d}m_{[G]}+\sum_{\pi\text{ cuspidal}}\sum_{\varphi\in\mathcal{B}(\pi)}\frac{\langle\phi,\varphi\rangle_{[G]}}{\|\varphi\|^{2}_{\pi}}\varphi(g)\\ +\int_{\Re(s)=0}\sum_{f\in\mathcal{B}(\mathcal{I}(s))}\frac{\langle\phi,\operatorname{Eis}(f)\rangle_{[G]}}{\|f\|^{2}_{\mathcal{I}(s)}}\operatorname{Eis}(f)(g)\frac{\,\mathrm{d}s}{4\pi i}.$
The right-hand side converges absolutely and uniformly on compacta. Moreover, in each summand the sum over $\mathcal{B}$ is invariant under a choice of orthogonal basis $\mathcal{B}$.
Lemma 3.5 is essentially given in [9, Lemma 2]. However, in [9] the test function $\phi$ is chosen to be inside $C^{\infty}([G])\cap L^{2}([G])$. Here we give a sketch to show that one may choose the test function from $C_{\mathrm{b}}^{2\ell}([G])$. The ideas of the proof are from [39, §2.6.5].
Proof.
We start with the $L^{2}$-version of the equality given in the statement of the lemma, which can be found in [14, Theorem, Chapter 1].
Let $\pi$ be cuspidal and $\varphi\in\mathcal{B}(\pi)$. Thus $\varphi$ is also an eigenvector of $\mathcal{D}$ with, say, eigenvalue $\lambda_{\varphi}$. Integrating by parts with respect to $\mathcal{D}$ for $l$ times we obtain
$\langle\phi,\varphi\rangle_{[G]}=\lambda_{\varphi}^{-l}\langle\mathcal{D}^{l}\phi,\varphi\rangle_{[G]}\ll\lambda_{\varphi}^{-l}\|\mathcal{D}^{l}\phi\|_{\infty}\|\varphi\|_{1}.$
Applying Lemma 3.4 we obtain the above is
$\frac{\langle\phi,\varphi\rangle_{[G]}}{\|\varphi\|^{2}_{\pi}}\ll\lambda_{\varphi}^{-l+d}\|\mathcal{D}^{l}\phi\|_{\infty}.$
Working similarly, for $f\in\mathcal{I}(s)$ we obtain
$\frac{\langle\phi,\operatorname{Eis}(f)\rangle_{[G]}}{\|f\|^{2}_{\mathcal{I}(s)}}\ll\lambda_{f}^{-l+d}\|\mathcal{D}^{l}\phi\|_{\infty}$
where $\lambda_{f}$ is the $\mathcal{D}$-eigenvalue of $f$ (note that $\operatorname{Eis}(f)$ is a $\mathcal{D}$-eigenfunction if and only if $f$ is). Once again applying Lemma 3.4 we have
$\varphi(g)\ll_{g}S_{d}(\varphi),\quad g\in[G].$
Thus the right-hand side of the claimed equation is bounded by
$\ll\|\phi\|_{1}+S_{\infty,l}(\phi)\left(\sum_{\pi\text{ cuspidal}}\sum_{\varphi\in\mathcal{B}(\pi)}\lambda_{\varphi}^{-l+2d}+\int_{\mathbb{R}}\sum_{f\in\mathcal{B}(\mathcal{I}(it))}\lambda_{f}^{-l+2d}\,\mathrm{d}t\right).$
The above $\varphi$-sum (resp. $f$-sum) can be realized as $\mathrm{trace}\left(\mathcal{D}^{-l+2d}|_{\pi}\right)$ (resp. $\mathrm{trace}\left(\mathcal{D}^{-l+2d}|_{\mathcal{I}(it)}\right)$) as $\mathcal{D}$ is positive definite. We make $l$ sufficiently large. Applying Lemma 3.1 for cuspidal $\pi$ and Weyl’s law (see, e.g., [39, §2.6.5 Corollary]) that
#$\#\{\pi\text{ cuspidal }\mid|\nu_{\pi}|\leq X\}\ll X^{O(1)},$
we conclude that the above $(\pi,\varphi)$-sum converges absolutely. Similarly, applying Lemma 3.1 for $\mathcal{I}(s)$ and noting that $\int_{\mathbb{R}}(1+|t|)^{-2}\,\mathrm{d}t<\infty$ we conclude that the above $t$-integral and $f$-sum converge absolutely.
Thus the right-hand side of the claimed equation converges absolutely, hence defines a continuous function of $g$. Thus the pointwise decomposition follows. ∎
4. Equidistribution with Average Fourier Decay
Lemma 4.1.
Let $\pi\ni\varphi$ be cuspidal or a unitary Eisenstein series, $x_{0}\in\mathbb{R}/\mathbb{Z}$, and $q\in\mathbb{N}$. Let $\mu$ be a Borel probability measure on $\mathbb{R}$. Then for any $\sigma>1$ there is a $d>0$ such that we have
$\int\varphi(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)=\sum_{0\neq|m|\leq|y|^{-\sigma}}\frac{\lambda_{\pi}(|m|)}{\sqrt{|m|}}W_{\varphi}(a(my))e(mx_{0})\widehat{\mu}(m/q)+O\left(|y|^{\frac{1}{2}}S_{2,d}(\varphi)\right)$
as $y\to 0$.
Proof.
We use the equation in Lemma 3.3 and integrate both sides in $x$ with respect to $\mu$. As $\mu$ is a probability measure and $\varphi_{0}$ is left $N$-invariant (as seen in Lemma 3.2) we obtain that
$\int\varphi(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)=\varphi_{0}(a(y))\\ +\sum_{0\neq|m|\leq|y|^{-\sigma}}\frac{\lambda_{\pi}(|m|)}{\sqrt{|m|}}W_{\varphi}(a(my))e(mx_{0})\widehat{\mu}(m/q)+O\left(|y|^{\frac{1}{2}}S_{2,d}(\varphi)\right).$
We conclude the proof after an application of the estimate in Lemma 3.2. ∎
Recall the definition of $\dim^{\star}_{\ell^{1}}\mu$ from (2.8).
Lemma 4.2.
Let $\mu$ be a Borel probability measure and $\ell<\dim^{\star}_{\ell^{1}}\mu$. Then for $q\in\mathbb{N}$ and $X\geq 1$ we have
$\sum_{|m|\leq X}|\widehat{\mu}(m/q)|\ll q^{\ell}X^{1-\ell}.$
Proof.
We may assume that $X\geq q$, otherwise, the claim is trivial. We see that
$\sum_{|m|\leq X}|\widehat{\mu}(m/q)|\leq 1+2\sum_{a=1}^{q}\sum_{n=0}^{[X/q]}|\widehat{\mu}(n+\tfrac{a}{q})|\ll q\sup_{0\leq\theta\leq 1}\sum_{n=0}^{[X/q]}|\widehat{\mu}(n+\theta)|.$
From (2.8), it follows that the above is
$\ll q[\tfrac{X}{q}]^{1-\ell}\leq q^{\ell}X^{1-\ell}$
as claimed. ∎
Lemma 4.3.
Let $x_{0}\in\mathbb{R}/\mathbb{Z}$ and $q\in\mathbb{N}$. Assume that there are an $\eta>0$ and $b,d>0$ such that for any $\pi\ni\varphi$ cuspidal or unitary Eisenstein series, we have
$\int\varphi(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)\ll q^{b}|y|^{\eta}S_{2,d}(\varphi)$
as $y\to 0$. Then there is a sufficiently large $\ell>0$ so that for any $\phi\in C_{\mathrm{b}}^{2\ell}([G])$ one has
$\int\phi(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)=\int_{[G]}\phi\,\mathrm{d}m_{[G]}+O\left(|y|^{\eta}q^{b}S_{\infty,\ell}(\phi)\right)$
as $y\to 0$.
Proof.
We start by spectrally decomposing $\phi$ as given in Lemma 3.5 at $g=n(x_{0}+\tfrac{x}{q})a(y)$ and integrate in $x$ with respect to $\mu$. We obtain
$\int\phi(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)-\int_{[G]}\phi=\sum_{\pi\text{ cuspidal}}\sum_{\varphi\in\mathcal{B}(\pi)}\frac{\langle\phi,\varphi\rangle_{[G]}}{\|\varphi\|^{2}_{\pi}}\int\varphi(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)\\ +\int_{\Re(s)=0}\sum_{f\in\mathcal{B}(\mathcal{I}(s))}\frac{\langle\phi,\operatorname{Eis}(f)\rangle_{[G]}}{\|f\|^{2}_{\mathcal{I}(s)}}\int\operatorname{Eis}(f)(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)\frac{\,\mathrm{d}s}{4\pi i}.$
The interchanges of above $\mu$ integral with spectral sum and integral are justified via absolute convergence of the spectral decomposition that is uniform in $x$ (which follows from Lemma 3.4) and finiteness of $\mu$. Now the proof follows exactly as the proof of Lemma 3.5 where we replace the estimates
$\varphi(g)\ll_{g}S_{2,d}(\varphi)\quad\text{ by }\quad\int\varphi(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)\ll q^{b}|y|^{\eta}S_{2,d}(\varphi),$
as given in the assumption. ∎
The following theorem is a stronger version of Theorem 2, under a slightly stronger assumption.
Proposition 4.1.
Let $x_{0}\in\mathbb{R}/\mathbb{Z}$, $q\in\mathbb{N}$ and $g_{0}:=n(x_{0})a(q^{-1})$. Let $\mu$ be a Borel probability measure such that
${\dim^{\star}_{\ell^{1}}}\mu>\frac{39}{64}=0.609375.$
Then there exist an $\eta>0$, $b>0$, and $\ell\in\mathbb{N}$ so that for any $\phi\in C_{\mathrm{b}}^{2\ell}(\Gamma\backslash G)$ we have
$\int\phi\left(g_{0}n(x)a(y)\right)\,\mathrm{d}\mu(x)=\int_{[G]}\phi\,\mathrm{d}m_{[G]}+O_{\mu}\left(q^{b}|y|^{\eta}S_{\infty,\ell}(\phi)\right),$
as $y\to 0$.
Proof.
First, note that
$g_{0}n(x)a(y)=n(x_{0}+\tfrac{x}{q})a(y)a(q^{-1}).$
Let $\pi\ni\varphi$ be a cuspidal representation or a unitary Eisenstein series. We start with the first term of the right-hand side of the equation in Lemma 4.1 with $\varphi$. Applying (3.5) with $7/64<\vartheta<\dim^{\star}_{\ell^{1}}\mu-\frac{1}{2}$, (3.7), and (3.13) with $A=-\frac{1}{2}+\vartheta$ we estimate
$\sum_{0\neq|m|\leq|y|^{-\sigma}}\frac{\lambda_{\pi}(|m|)}{\sqrt{|m|}}W_{\varphi}(a(my))e(mx_{0})\widehat{\mu}(m/q)$ $\ll S_{2,d}(\varphi)\sum_{|m|\leq|y|^{-\sigma}}|m|^{\vartheta-\frac{1}{2}}|my|^{\frac{1}{2}-\vartheta}|\widehat{\mu}(m/q)|$
$\ll S_{2,d}(\varphi)|y|^{\frac{1}{2}-\vartheta}\sum_{|m|\leq|y|^{-\sigma}}|\widehat{\mu}(m/q)|$
for some $d>0$ independent of $\varphi$ and $q$.
Fix $\frac{1}{2}+\vartheta<b<\dim^{\star}_{\ell^{1}}\mu$. Then from Lemma 4.2 we obtain the last sum above is $O\left(q^{b}|y|^{-\sigma(1-b)}\right)$. Hence, from Lemma 4.1 we obtain
$\int\varphi(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)\ll S_{2,d}(\varphi)\left(|y|^{\frac{1}{2}}+q^{b}|y|^{b-(\frac{1}{2}+\vartheta)-(\sigma-1)(1-b)}\right).$
Choosing $10<\sigma-$ 1 sufficiently small we obtain that
$\int\varphi(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)\ll q^{b}|y|^{\eta}S_{2,d}(\varphi)$
for some $\eta>0$. We apply Lemma 4.3 with $\phi_{q}:=\phi(\cdot a(q^{-1}))$ to obtain
$\int\phi(g_{0}n(x)a(y))\,\mathrm{d}\mu(x)=\int\phi_{q}(n(x_{0}+\tfrac{x}{q})a(y))\,\mathrm{d}\mu(x)=\int_{[G]}\phi_{q}\,\mathrm{d}m_{[G]}+O\left(|y|^{\eta}q^{b}S_{\infty,\ell}(\phi_{q})\right).$
Finally, noting that (see, e.g., [27, Lemma 3.3 (1)])
$S_{\infty,\ell}(\phi_{q})\ll\|\mathrm{Ad}(a(q^{-1}))\|^{\ell}S_{\infty,\ell}(\phi)\ll q^{O(1)}S_{\infty,\ell}(\phi)$
and by measure invariance
$\int_{[G]}\phi_{q}=\int_{[G]}\phi,$
we conclude the proof. ∎
Proof of Theorem 1.
The proof follows similarly to the proof of Proposition 4.1, but with $x_{0}=0$ and $q=1$, and $\dim^{\star}_{\ell^{1}}$ replaced by $\dim_{\ell^{1}}$. ∎
Proof of Corollary 1.
Note that for any two Borel probability measures $\mu_{1}$ and $\mu_{2}$ we have
$\widehat{\mu_{1}\ast\mu_{2}}(\xi)=\widehat{\mu}_{1}(\xi)\widehat{\mu}_{2}(\xi).$ (4.1)
Hence by Cauchy–Schwarz we have
$\sum_{|m|\leq X}|\widehat{\mu_{1}\ast\mu_{2}}(m)|\leq\left(\sum_{|m|\leq X}|\widehat{\mu}_{1}(m)|^{2}\right)^{1/2}\left(\sum_{|m|\leq X}\widehat{\mu}_{1}(m)|^{2}\right)^{1/2}\ll X^{1-\frac{l_{1}+l_{2}}{2}}$
[…]
$\dim_{\ell^{1}}\mu\ast\mu\geq\frac{s_{1}+s_{2}}{2}.$
The corollary now follows from Theorem 1. ∎
Proof of Corollary 2.
Applying (4.1) and the fact that $|\widehat{\mu}_{j}(\xi)|\leq 1$ for all $\xi$ we conclude that $\dim_{\ell^{1}}\mu\ast\nu\geq\dim_{\ell^{1}}\mu$. ∎
Now we prove Theorem 2.
Proof of Theorem 2.
From [12, Theorem 2.6] we recall that for $D$ in arithmetic progression with #$\#D=l$ one has
$\dim_{\ell^{1}}\mu_{b,D}\geq\frac{\log l}{\log b}-\frac{\log(4+\log(2l))}{\log b}.$ (4.2)
Suppose $\tfrac{39}{64}<s<1$. Let $b(s)\in\mathbb{N}$ such that for every $b\geq b(s)$ we have
$s-\frac{\log(4+\log(2b))}{\log b}>\frac{39}{64}\quad\text{and}\quad b-b^{s}\geq 2.$ (4.3)
Now fix $b\geq b(s)$ and let $2\leq l_{1}\in(b^{s},b)\cap\mathbb{N}$. Let $K_{b,D}$ be such that $D$ is any subset of $\{0,\dots,b-1\}$ with $l_{1}$ elements and they are in an arithmetic progression. $K_{b,D}$. Thus
$\dim_{\mathcal{H}}K_{b,D}=\frac{\log l_{1}}{\log b}=:s_{1}\geq s.$
Note that
$s_{1}-\frac{\log(4+\log(2l_{1}))}{\log b}\geq s_{1}-s+s-\frac{\log(4+\log(2b))}{\log b}>\frac{39}{64}.$
Thus from (4.2) we have $\dim_{\ell^{1}}\mu_{b,D}>\tfrac{39}{64}$, consequently, the proof follows from Theorem 1. ∎
Proof of Theorem 3.
The proof immediately follows from (2.3), discussion around (1.3), and proof of Theorem 2. ∎
Lemma 4.4.
Let $\mu$ be a self-similar measure whose underlying attractor is a missing digit Cantor set. Then
$\dim_{\ell^{1}}\mu=\dim^{\star}_{\ell^{1}}\mu.$
Moreover, the same is true when the attractor is a shifted missing digit Cantor set.
Proof.
We first prove when $\mu$ is a self-similar measure on $K_{b,D}$. The proof of this is essentially available in [12], although was not pointed out. Note that using (2.9) and [12, Theorem 4.2], the lemma follows upon showing that for all $L\in\mathbb{N},$
$f_{L}:=\frac{-\log(\max_{x}b^{-L}\sum_{i=0}^{b^{L}-1}S_{L}(x+\frac{i}{b^{L}}))}{\log b^{L}}\leq\dim^{\star}_{\ell^{1}}\mu,$ (4.4)
where $S_{L}$ is as defined in [12, Equation 4.3]. Then by the argument as in [12, §4.1], it is enough to prove (4.4) for $L=1.$ Now in the proof of [12, Lemma 4.3] authors show that for any $N\in\mathbb{Z}_{\geq 0}$ and $\theta\in\mathbb{R},$
$\sum_{m=0}^{b^{N}-1}|\widehat{\mu}(m+\theta)|\leq b^{N(1-f_{1})}.$
Hence
$\sup_{0\leq\theta\leq 1}\sum_{m=0}^{b^{N}-1}|\widehat{\mu}(m+\theta)|\leq b^{N(1-f_{1})},$
which implies (4.4).
To see the second part, we note if $\mu$ is a self-similar measure on $K_{b,D}+x_{0}$ for some $x_{0}\in\mathbb{R}$ then by uniqueness of the self-similar measures $\mu=\mu_{1}\ast\delta_{x_{0}}$ where $\mu_{1}$ is a self-similar measure on $K_{b,D}$; see (2.3). Thus $|\widehat{\mu}|=|\widehat{\mu}_{1}|$. ∎
Proposition 4.2.
Let $\mu$ be a self-similar measure with the IFS being irreducible and having OSC. If
$\dim^{\star}_{\ell^{1}}\mu>0.609375,$
then for any $\psi$ non-increasing monotonic function,
$\mu(\mathcal{W}(\psi))=1\quad\text{ if }\quad\sum\psi(q)=\infty.$
Proof.
The proposition follows combining Proposition 4.1 and [27, Theorem 12.1]. ∎
Proof of Theorem 6.
Let $\mu$ be a self-similar measure whose underlying attractor is a shifted missing digit Cantor set. Therefore the corresponding IFS satisfies OSC and is irreducible. Also by Lemma 4.4, $\dim_{\ell^{1}}\mu=\dim^{\star}_{\ell^{1}}\mu$. Thus the hypothesis of Proposition 4.1 is valid. Hence the proof follows from Propostion 4.2. ∎
Proof of Theorem 4.
Let $\rho$ be sufficiently small such that $(\rho^{-0.609375},\rho^{-1}]\cap\mathbb{Z}_{\geq 2}\neq\varnothing$. Choose $l$ from the above interval. Let $\mu$ be the self-similar measure associated to the IFS $\{\rho(x+i),i=0,\cdots,l-1\}$ and the uniform probability vector. Since this IFS satisfies OSC, $\mu$ is $\frac{-\log l}{\log\rho}$-AD regular (see §2.1). Note that $\mu\ast\mu$ is a self-similar measure with (equal) contraction $\rho$; see [50, pp.809]. By (2.7), $\dim_{\ell^{1}}\mu\ast\mu=\dim_{\ell^{2}}\mu=\frac{-\log l}{\log\rho}>0.609375,$ therefore the corollary follows from Theorem 1. ∎
Remark 4.1.
Note that in the proof of Theorem 4, when $\rho^{-1}$ is the (positive) $n$’th root of a positive integer then $\mu\ast\mu$ is not absolutely continuous, since $\widehat{\mu}\nrightarrow 0$ in this case, which follows from (2.5).
5. Equidistribution with Pointiwse Fourier Decay
Lemma 5.1.
There is a $d>0$ such that for every $\delta>0$ and $0<\eta<\min(\tfrac{1}{2},\delta)$ we have the following. Let $\pi$ be either cuspidal representation or unitary Eisenstein series and $\varphi\in\pi$. Then we have
$\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{|m|^{1+\delta}}W_{\varphi}(a(my))\ll_{\eta}|y|^{\eta}S_{2,d}(\varphi)$
as $y\to 0$.
Proof.
We use the the bound of Whittaker function (3.13) with $A=-\eta>-\min(\tfrac{1}{2},\delta)$. Then we bound the sum in the lemma by
$\ll|y|^{\eta}S_{2,d}(W_{\varphi})\sum_{m\geq 1}\frac{|\lambda_{\pi}(m)|}{m^{1+\delta-\eta}}.$
Thus it suffices to show that the above sum is $\nu_{\pi}^{O(1)}$. To see this, first note that for any $X>1$ applying summation by parts we have
$\sum_{m\leq X}\frac{|\lambda_{\pi}(m)|}{m^{1+\delta-\eta}}=X^{-(1+\delta-\eta)}\sum_{m\leq X}|\lambda_{\pi}(m)|+\int_{1}^{X}\frac{\sum_{m\leq t}|\lambda_{\pi}(m)|}{t^{2+\delta-\eta}}\,\mathrm{d}t.$
Applying Cauchy–Schwarz and employing (3.8) it follows that
$\sum_{m\leq t}|\lambda_{\pi}(m)|\ll_{\epsilon}t^{1+\epsilon}\nu_{\pi}^{\epsilon}.$
As $11+\delta-\eta>$ 1, the proof follows after letting $X\to\infty$. ∎
Now we will estimate a similar sum as in Lemma 5.1 but $1+\delta$ is replaced by $\tfrac{1}{2}+\delta$, namely,
$\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{|m|^{\frac{1}{2}+\delta}}e(m\alpha)W_{\varphi}(a(my)),\quad\alpha\in\mathbb{R}.$
To prove a polynomial decay in $y$ for the above sum we can not mimic the proof of Lemma 5.1, as in this proof the absolute convergence of the Dirichlet series $\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{|m|^{s}}$ for $\Re(s)>1$ is crucially used, which will not be available in this case.
We will analyse this sum using a Voronoi-type argument. This argument rather uses the meromorphic properties of the Dirichlet series of the additive twisted Hecke eigenvalues. Such an argument is common in literature when $\pi$ is cuspidal. However, we are unable to find a reference for the same when $\pi$ is Eisenstein. Below we give a proof for the required estimate for all unitary automorphic $\pi$.
As a preparation, we first approximate $\alpha$ by a nonzero rational number $\frac{p}{q}$ with $q\in\mathbb{N}$ in its reduced form so that
$|\xi|\leq\frac{1}{\sqrt{|y|}q},\quad\xi:=\frac{1}{|y|}\left(\alpha-\frac{p}{q}\right),\quad 1\leq q\leq|y|^{-\frac{1}{2}}.$ (5.1)
The above is guaranteed because of the Dirichlet’s approximation. Finally, we abbreviate $W_{\varphi}(\cdot n(\xi))$ by $W^{\xi}$. We start with the following Mellin-theoretic property of $W^{\xi}$.
Lemma 5.2.
If $z$ is away from any pole of $Z(\cdot,W^{\xi})$ then
$Z(z,\varepsilon,W^{\xi})\ll_{N}(1+|z|)^{2\min(0,\Re(z))}\left(\frac{1+|\xi|}{1+|z|}\right)^{N}S_{2,d}(W_{\varphi})$
for some $d>0$ depending only on $N$.
Proof.
If $\Re(z)\geq 0$ then
$Z(z,\varepsilon,W^{\xi})=\int_{\mathbb{R}^{\times}}W(a(t))e(t\xi)|t|^{z}\operatorname{sgn}(t)^{\varepsilon}\,\mathrm{d}^{\times}t$
converges absolutely. We integrate by parts the above with $(t\partial_{t})^{N}$ and apply (3.13) to obtain
$Z(z,\varepsilon,W^{\xi})\ll_{N}\left(\frac{1+|\xi|}{1+|z|}\right)^{N}S_{2,d}(W)$
If $\Re(z)<0$ and $z$ is not a pole of $Z(W^{\xi})$ then applying local functional equation as in (3.14) we obtain
$Z(z,\varepsilon,W^{\xi})=\gamma(\tfrac{1}{2}-z,\pi)Z(-z,\varepsilon,W^{\xi}).$
We bound the gamma factor by
$\gamma(\tfrac{1}{2}-z,\pi)\ll(1+|z|)^{2\Re(z)}\nu_{\pi}^{O(1)}$
and apply the above estimate for $Z(-z,W^{\xi})$ for $-\Re(z)>0$. ∎
Lemma 5.3.
There is a $d>0$ such that for every $0<\delta\leq\tfrac{1}{2}$ and every $0<\eta<\tfrac{\delta}{2}$ we have the following. Let $\pi$ be a cuspidal representation of $G$ and $\varphi\in\pi$ be an automorphic form. Let $p,q,\xi$ be as in (5.1). Then
$\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{|m|^{\frac{1}{2}+\delta}}e\left(m\tfrac{p}{q}\right)W^{\xi}(a(my))\ll_{\eta}|y|^{\eta}S_{2,d}(\varphi)$
as $y\to 0$.
Proof.
We start with the Mellin expansion of $W^{\xi}$ given by
$W^{\xi}(a(y))=\frac{1}{2}\sum_{\varepsilon\in\{0,1\}}\operatorname{sgn}(y)^{\varepsilon}\int_{\Re(z)=\sigma}Z(z,\varepsilon,W^{\xi})|y|^{-z}\frac{\,\mathrm{d}z}{2\pi i}$
for some sufficiently positive $\sigma$, where $Z(\cdots)$ is as in (3.12). Using this, we write the sum in the lemma as
$\frac{1}{2}\sum_{\varepsilon\in\{0,1\}}\operatorname{sgn}(y)^{\varepsilon}\int_{\Re(z)=\sigma}Z(z,\varepsilon,W^{\xi})|y|^{-z}\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)\operatorname{sgn}(m)^{\varepsilon}e\left(m\tfrac{p}{q}\right)}{|m|^{\frac{1}{2}+z+\delta}}\frac{\,\mathrm{d}z}{2\pi i}.$
The above interchange of the $m$-sum and the $z$-integral is justified because the $m$-sum converges absolutely for sufficiently large $\sigma$ (which follows from (3.5) and (3.7)) and $Z(z,\varepsilon,W^{\xi})$ decays rapidly in $z$ (which follows from Lemma 5.2).
Let us consider the $\varepsilon=0$ case to ease the notations; the $\varepsilon=1$ case will be similar. Also, from now on, we will drop $\varepsilon(=0)$ from the notations. Moreover, we only consider the $m>0$ part of the above sum; the $m<0$ part can be treated similarly by replacing $\tfrac{p}{q}$ by its negative.
Correspondingly, we can rewrite the $\varepsilon=0$ summand in the above expression as
$\int_{\Re(z)=\sigma}Z(z,W^{\xi})|y|^{-z}L(\tfrac{1}{2}+z+\delta,\tfrac{p}{q},\pi)\frac{\,\mathrm{d}z}{2\pi i}.$ (5.2)
where we define
$L\left(z,\tfrac{p}{q},\pi\right):=\sum_{m=1}^{\infty}\frac{\lambda_{\pi}(m)e\left(m\tfrac{p}{q}\right)}{m^{z}}.$
It follows from (3.8) that the above converges absolutely for $\Re(z)>1$.
As $\pi$ is cuspidal, it is known that $L\left(z,\tfrac{p}{q},\pi\right)$ is entire in $z$; see [33, Appendix 3-4]. First, we will bound the $L\left(z,\tfrac{p}{q},\pi\right)$ for $0\leq\Re(z)\leq 1$. Note that using (3.8) we can bound
$L\left(z,\tfrac{p}{q},\pi\right)\ll_{\Re(z)}1,\quad\Re(z)>1.$
On the other hand, $L\left(z,\tfrac{p}{q},\pi\right)$ satisfies the functional equation [33, eq.(A.10),(A.12),(A.13)] of the form
$L\left(z,\tfrac{p}{q},\pi\right)=q^{1-2z}\left[L\left(1-z,-\tfrac{\overline{p}}{q},{\pi}\right)\gamma_{1}(z,\pi)+L\left(1-z,\tfrac{\overline{p}}{q},{\pi}\right)\gamma_{2}(z,\pi)\right]$
for some complex meromorphic functions $\gamma_{j}(z,\pi)$ (given in terms of $C^{\pm}$ as in [33]) satisfying
$\gamma_{j}(z,\pi)\ll_{\Re(z)}(1+|z|)^{1-2\Re(z)}\nu_{\pi}^{O_{\Re(z)}(1)}$
as long as $z$ is a fixed distance away from any pole of $\gamma_{j}(\cdot,\pi)$. The above bound follows from a standard application of Stirling’s estimate of $\Gamma$-function; e.g., see [14, Proposition 9.5-9.6]. Hence, by the functional equation of the $L$-function and the above bounds of the $\gamma$-factor we have
$L\left(z,\tfrac{p}{q},\pi\right)\ll_{\Re(z)}\left(q(1+|z|)\right)^{1-2\Re(z)}\nu_{\pi}^{O_{\Re(z)}(1)},\quad\Re(z)<0.$
Thus applying Phragmén–Lindelöf convexity principle, we conclude
$L\left(z,\tfrac{p}{q},\pi\right)\ll_{\epsilon}(q(1+|z|))^{1-\Re(z)+\epsilon}\nu_{\pi}^{O(1)},\quad 0\leq\Re(z)\leq 1.$ (5.3)
Now using the holomorphic and decay properties of $Z(z,W^{\xi})$ for $\Re(z)>-\frac{1}{2}$ as in Lemma 5.2, we shift the contour of (5.2) to $\Re(z)=\sigma$ where $-\frac{1}{2}<\sigma<0$ without crossing any poles of the integrand. Applying (5.3) and Lemma 5.2 we estimate the shifted contour integral (5.2) by
$\ll_{\sigma,N}\nu_{\pi}^{O(1)}|y|^{-\sigma}q^{\frac{1}{2}-\sigma-\delta}\int_{\mathbb{R}}(1+|z|)^{\frac{1}{2}+\sigma-\delta}\left(\frac{1+|\xi|}{1+|z|}\right)^{N}\,\mathrm{d}z\ll_{\sigma}|y|^{-\sigma}q^{\frac{1}{2}-\sigma-\delta}(1+|\xi|)^{\frac{3}{2}+\sigma-\delta}\nu_{\pi}^{O(1)}.$
We choose $\sigma=-\tfrac{1}{2}+\epsilon$ and apply (5.1) to estimate the above by $O\left(|y|^{\frac{\delta}{2}-\epsilon}\nu_{\pi}^{O(1)}\right)$. ∎
Lemma 5.4.
There is an $\ell>0$ such that for every $0<\delta\leq\tfrac{1}{2}$, and every $0<\eta<\tfrac{\delta}{2}$ we have the following. For any $10\leq\alpha<$ 1 and any $\phi\in C^{2\ell}_{\mathrm{b}}([G])$ we have
$\sum_{\pi\text{ cuspidal}}\sum_{\varphi\in\mathcal{B}(\pi)}\frac{\langle\phi,\varphi\rangle_{[G]}}{\|\varphi\|^{2}_{\pi}}\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{|m|^{\frac{1}{2}+\delta}}W_{\varphi}(a(my))e(m\alpha)\ll_{\eta}|y|^{\eta}S_{\infty,\ell}(\phi)$
as $y\to 0$.
Proof.
Using unipotent equivariance of $W$ as in (3.11) and Dirichlet approximation (5.1) we deduce
$W_{\varphi}(a(my))e(m\alpha)=W^{\xi}(a(my))e\left(m\tfrac{p}{q}\right).$
Thus we write the sum in the Lemma as
$\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{|m|^{\frac{1}{2}+\delta}}W^{\xi}(a(my))e\left(m\tfrac{p}{q}\right).$
We conclude using Lemma 5.3 that
$\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{|m|^{\frac{1}{2}+\delta}}W_{\varphi}(a(my))e(m\alpha)\ll|y|^{\eta}S_{2,d}(\varphi).$
for cuspidal $\pi$. Working as in the proof of Lemma 3.5 we see that
$\sum_{\pi\text{ cuspidal}}\sum_{\varphi\in\mathcal{B}(\pi)}\frac{\langle\phi,\varphi\rangle_{[G]}}{\|\varphi\|^{2}_{\pi}}S_{2,d}(\varphi)$
converges absolutely and is $O\left(S_{\infty,\ell}(\phi)\right)$. ∎
Lemma 5.5.
There is a $d>0$ such that for every $0<\delta\leq\tfrac{1}{2}$ and every $0<\eta<\tfrac{\delta}{2}$ we have the following. Let $\pi$ be a unitary Eisenstein series with parameter $s\in i\mathbb{R}$ and $\varphi=\operatorname{Eis}(f)$ with $f\in\mathcal{I}(s)$. Let $p,q,\xi$ be as in (5.1). Then
$\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{|m|^{\frac{1}{2}+\delta}}e\left(m\tfrac{p}{q}\right)W^{\xi}(a(my))=\mathcal{M}+O_{\eta}\left(|y|^{\eta}S_{2,d}(\varphi)\right)$
where
:=±$\mathcal{M}:=\sum_{\pm}Z(\tfrac{1}{2}-\delta\pm s,W^{\xi})|y|^{\delta}\left(\sqrt{|y|}q\right)^{-1\mp 2s}\frac{\zeta(1\pm 2s)}{\zeta(1+2s)},$ (5.4)
as $y\to 0$.
Proof.
Working as in the proof of cuspidal case (with the same conventions and notations), we write the sum in the lemma as
$\int_{\Re(z)=\sigma}|y|^{-z}Z(z,W^{\xi})L(\tfrac{1}{2}+z+\delta,\tfrac{p}{q},\pi)\frac{\,\mathrm{d}z}{2\pi i}$ (5.5)
for some $\sigma>\tfrac{1}{2}-\delta$. Here from (3.6) it follows that
$L\left(z,\tfrac{p}{q},\pi\right)=\frac{D\left(z+s,2s,\tfrac{p}{q}\right)}{\zeta(1+2s)}$
where
$D\left(z,s,\tfrac{p}{q}\right):=\sum_{m=1}^{\infty}\frac{\tau_{s}(m)e\left(m\tfrac{p}{q}\right)}{m^{z}},$
as defined in [6, pp.228]. As before, we first determine holomorphic and growth properties of $L\left(z,\tfrac{p}{q},\pi\right)$ for $0\leq\Re(z)\leq 1$ and $\Re(s)=0$.
Clearly, for $\Re(z)>1$ we have $D\left(z+s,2s,\tfrac{p}{q}\right)\ll_{\Re(z)}1$. On the other hand, $D$ satisfies a functional equation, namely,
$D\left(z+s,2s,\tfrac{p}{q}\right)=-\frac{2}{q}\left(\frac{q}{2\pi}\right)^{2-2z}\Gamma(1-z+s)\Gamma(1-z-s)\\ \times\left[\cos(\pi z)D(1-z-s,-2s,-\tfrac{\overline{p}}{q})-\cos(\pi s)D(1-z-s,-2s,\tfrac{\overline{p}}{q})\right],$
which follows from [6, Lemma 2]. Using Stirling’s estimate (see, e.g., [14, eq.(9.1)])
$\Gamma(\sigma+it)\ll_{\sigma}\exp\left(-\frac{\pi}{2}|t|\right)|\sigma+it|^{\sigma-\frac{1}{2}},\quad\sigma,t\in\mathbb{R},\sigma\notin\mathbb{Z}_{\leq 0}$
for $\Re(s)=0$ we estimate
$\Gamma(1-z+s)\Gamma(1-z-s)\ll_{\Re(z)}\exp(-\pi\max(|\Im(z)|,|s|))(1+|z|^{2}+|s|^{2})^{\frac{1}{2}-\Re(z)}.$
Moreover, using that $\cos(\pi z)\asymp_{\Re(z)}\exp(\pi|\Im(z)|)$, we conclude
$D\left(z+s,2s,\tfrac{p}{q}\right)\ll(q(1+|z|))^{1-2\Re(z)}(1+|s|)^{O_{\Re(z)}(1)},\quad\Re(z)<0,\Re(s)=0.$
Thus, using Phragmén–Lindelöf convexity principle and $\zeta(1+s)^{-1}\ll(1+|s|)^{\epsilon}$ for $\Re(s)=0$ (see [53, pp.50-51]) we obtain
$L\left(z,\tfrac{p}{q},\pi\right)\ll_{\epsilon}(q(1+|z|))^{1-\Re(z)+\epsilon}\nu_{\pi}^{O(1)},\quad 0\leq\Re(z)\leq 1.$ (5.6)
On the other hand, from [6, Lemma 2] we deduce that $L\left(z,\tfrac{p}{q},\pi\right)$ has poles in $z$ exactly at the poles of
$q^{1-2z}\frac{\zeta(z-s)\zeta(z+s)}{\zeta(1+2s)}$
which are at $z=1\pm s$ and of order $1$ with residues $q^{-1\mp 2s}\frac{\zeta(1\pm 2s)}{\zeta(1+2s)}$, respectively.
Now, as in the proof of the cuspidal case, using holomorphic and decay properties of $Z(z,W^{\xi})$ for $\Re(z)>-\frac{1}{2}$ as in Lemma 5.2 we shift the contour of (5.5) to $\Re(z)=\sigma$ where $-\frac{1}{2}<\sigma<0$. Using Cauchy’s theorem we can write (5.5) as the sum of
=±$\mathcal{M}=\sum_{\pm}Z(\tfrac{1}{2}-\delta\pm s,W^{\xi})|y|^{\delta}\left(\sqrt{|y|}q\right)^{-1\mp 2s}\frac{\zeta(1\pm 2s)}{\zeta(1+2s)},$
and the shifted contour integral (5.5). Working exactly as in the cuspidal case, in particular, using Lemma 5.2 and (5.6), and choosing $\sigma=-\frac{1}{2}+\epsilon$ we bound the shifted contour integral (5.5) by $O_{\epsilon}\left(|y|^{\frac{\delta}{2}-\epsilon}\nu_{\pi}^{O(1)}\right)$. ∎
Lemma 5.6.
There is an $\ell>0$ such that for every $0<\delta\leq\tfrac{1}{2}$ and every $0<\eta<\tfrac{\delta}{2}$ we have the following. For any $10\leq\alpha<$ 1 and any $\phi\in C^{2\ell}_{\mathrm{b}}([G])$ we have
$\int_{\Re(s)=0}\sum_{f\in\mathcal{B}(\mathcal{I}(s))}\frac{\langle\phi,\operatorname{Eis}(f)\rangle_{[G]}}{\|f\|^{2}_{\mathcal{I}(s)}}\sum_{m\neq 0}\frac{\lambda_{s}(m)}{|m|^{\frac{1}{2}+\delta}}W_{f}(a(my))e(m\alpha)\frac{\,\mathrm{d}s}{2\pi i}\ll_{\eta}|y|^{\eta}S_{\infty,\ell}(\phi)$
as $y\to 0$. Here $\lambda_{s}$ we mean $\lambda_{\pi}$ where $\pi$ is the Eisenstein series with parameter $s$.
Proof.
We work as in the proof of Lemma 5.4 to reduce to $\alpha=\tfrac{p}{q}$ and $W_{f}$ to $W^{\xi}$ where $\xi,p,q$ are as in (5.1). Then we apply Lemma 5.5 for $\pi$ Eisenstein to write the sum in the lemma as
$|y|^{\delta}\int_{\Re(s)=0}\left(\sqrt{|y|}q\right)^{-1-2s}H(\tfrac{1}{2}-\delta+s)\frac{\,\mathrm{d}s}{2\pi i}\\ +|y|^{\delta}\int_{\Re(s)=0}\left(\sqrt{|y|}q\right)^{-1+2s}\frac{\zeta(1-2s)}{\zeta(1+2s)}H(\tfrac{1}{2}-\delta-s)\frac{\,\mathrm{d}s}{2\pi i}\\ +O\left(|y|^{\eta}\int_{\mathbb{R}}\sum_{f\in\mathcal{B}(\mathcal{I}(it))}\frac{|\langle\phi,\operatorname{Eis}(f)\rangle_{[G]}|}{\|f\|^{2}_{\mathcal{I}(it)}}S_{2,d}(\operatorname{Eis}(f))\,\mathrm{d}t\right),$ (5.7)
where
$H(z)=H(z;W^{\xi},\phi,s):=\sum_{f\in\mathcal{B}(\mathcal{I}(s))}\frac{\langle\phi,\operatorname{Eis}(f)\rangle_{[G]}}{\|f\|^{2}_{\mathcal{I}(s)}}Z(z,W^{\xi}).$
Working as in the proof of Lemma 3.5 we see that the integral-sum in the third summand in (5.7) is convergent and consequently, the third summand is $O\left(|y|^{\eta}S_{\infty,\ell}(\phi)\right)$.
We now deal with the second summand in (5.7). Using (3.10) and (3.16) we write
$\frac{\zeta(1-2s)}{\zeta(1+2s)}H(\tfrac{1}{2}-\delta-s)$ $=\sum_{f\in\mathcal{B}(\mathcal{I}(s))}\frac{\langle\phi,\operatorname{Eis}(M^{\ast}(s)f)\rangle_{[G]}}{\|M^{\ast}(s)f\|^{2}_{\mathcal{I}(-s)}}Z(\tfrac{1}{2}-\delta-s,W^{\xi})$
$=\sum_{f\in\mathcal{B}(\mathcal{I}(-s))}\frac{\langle\phi,\operatorname{Eis}(f)\rangle_{[G]}}{\|f\|^{2}_{\mathcal{I}(-s)}}Z(\tfrac{1}{2}-\delta-s,W^{\xi})$
The last equality follows due to (3.9) and noting that $\{M^{\ast}(s)f\}_{f\in\mathcal{B}(\mathcal{I}(s)}$ forms an orthogonal basis of $\mathcal{I}(-s)$. Using the above and changing variable $s\mapsto-s$ we see the second summand in (5.7) equals
$|y|^{\delta}\int_{\Re(s)=0}\left(\sqrt{|y|}q\right)^{-1-2s}H(\tfrac{1}{2}-\delta+s)\frac{\,\mathrm{d}s}{2\pi i}$
which is the same as the first summand in (5.7) on which we focus next.
We first analytically continue $H(z,..,s)$ in neighbourhood of $\Re(s)=0$. Note that $\overline{\operatorname{Eis}(f)}=\operatorname{Eis}(\bar{f})$ and $\bar{f}\in\mathcal{I}(-s)$ for $\Re(s)=0$. Thus changing $s$ to $-s$ we give a holomorphic realization of $\langle\phi,\operatorname{Eis}(f)\rangle_{[G]}$. Recalling that $\|f\|_{\mathcal{I}(s)}$ is $s$-independent and meromorphic properties of $\operatorname{Eis}(f)$ we obtain that $\frac{\langle\phi,\operatorname{Eis}(f)\rangle_{[G]}}{\|f\|^{2}_{\mathcal{I}(s)}}$ is holomorphic in $-\tfrac{1}{2}<\Re(s)\leq 0$. On the other hand, recalling holomorphic properties of $Z$ and $W$ we get that $Z(\tfrac{1}{2}-\delta+s)$ is holomorphic in $-\tfrac{1}{2}<\Re(s)\leq 0$. Finally, incorporating Remark 3.1, using the bound in Lemma 5.2, and following the proof of Lemma 3.5 we obtain that the sum defining $H(\tfrac{1}{2}-\delta+s)$ converges absolutely in $-\tfrac{1}{2}<\Re(s)\leq 0$ and thus defines a holomorphic function in the same region. Moreover, the same proof yields that
$H(\tfrac{1}{2}-\delta+s)\ll_{N,\Re(s)}(1+|s|)^{-N}\|\mathcal{D}^{\ell}\phi\|_{\infty},\quad 0\leq\Re(s)<\tfrac{1}{2},\quad-\tfrac{1}{2}<\Re(s)\leq 0$
for some $\ell>0$ depending on $N$.
Now we shift the contour of the first summand in (5.7) to $\Re(s)=-\tfrac{1}{2}+\epsilon$ without crossing any pole. Using the last estimate of $H$ with sufficiently large $N$ we bound the shifted integral by $\ll|y|^{\delta-\epsilon}\|\mathcal{D}^{\ell}\phi\|_{\infty}$ completing the proof. ∎
Proof of Theorem 5.
Without loss of generality we assume $K=1$ and drop the subscripts from $\delta_{1}\in(0,\tfrac{1}{2}],\alpha_{1}\in\mathbb{R},\beta_{1}\in\mathbb{C}$. Let $\pi$ be either cuspidal or unitary Eisenstein series and $\varphi\in\pi$. Applying absolute convergent Fourier expansion of $\varphi-\varphi_{0}$, as in (3.4), and bound of $\varphi_{0}(a(y))$, as in Lemma 3.2 we write
$\int\varphi(n(x)a(y))\,\mathrm{d}\mu(x)=\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|}{\sqrt{|m|}}W_{\varphi}(a(my))\widehat{\mu}(m)+O\left(|y|^{\frac{1}{2}}S_{2,d}(\varphi)\right).$
Now employing the expression of $\widehat{\mu}$ as in Theorem 5 and applying Lemma 5.1 we deduce that the right hand side above is
$\beta\sum_{m\neq 0}\frac{\lambda_{\pi}(|m|)}{|m|^{\frac{1}{2}+\delta}}W_{\varphi}(a(my))e(m\alpha)+O\left(|y|^{\eta}S_{2,d}(\varphi)\right).$
Starting with the spectral decomposition as in Lemma 3.5 then applying Lemma 5.4 and Lemma 5.6, and working as in the proof of Lemma 4.3, we conclude the proof. ∎
5.1. Stationary Phase
In this subsection, we prove Corollary 4. We start with the stationary phase estimate of the Fourier transform of $\mu^{w,f}$ as in Corollary 4.
Proposition 5.1.
Let $f,w$ and $\mu^{w,f}$ be as in Corollary 4. Let $\{x_{i}\}_{i=1}^{n}$ be the set of points in the support of $w$ and let $\{k_{i}\}_{i=1}^{n}$ be the multi-set of positive integers so that $f^{\prime}$ has a zero of order $k_{i}-1$ at $x_{i}$.
Then there exist $\{a_{i,j}\}_{1\leq i\leq n}^{j\in\mathbb{N}}$ only depending on $f$ and $w$ so that for any $N>0$
$\widehat{\mu^{w,f}}(\xi)=\sum_{i=1}^{n}e\left(\xi f(x_{i})\right)\sum_{j=0}^{N-1}a_{i,j}\xi^{-\frac{j+1}{k_{i}}}+O_{f,w,N}\left(\left(1+|\xi|\right)^{-\frac{N+1}{\max_{i}{k_{i}}}}\right),$
as $|\xi|\to\infty$.
Proof.
We first discuss how the above proposition can be reduced to a simpler one. Modifying the support of $w$ and using [49, Proposition 1], without loss of generality, we can assume $i=1$, i.e., there is a unique stationary point of $f$ in the support of $w$ (equivalently, $x_{1}=x_{2}=\cdots=x_{n}$). Also by changing $f$ with $x\mapsto f(x+x_{1})-f(x_{1})$, we can also assume without loss of generality, that $x_{1}=0$ and
$f(0)=f^{\prime}(0)=\cdots=f^{(k-1)}(0)=0,\quad f^{(k)}(0)\neq 0.$
To prove Proposition 5.1 it is enough to show that for the above $f$ for any $N>0$ we have
$\widehat{\mu^{w,f}}(\xi)=\sum_{j=0}^{N-1}a_{j}\xi^{-\frac{j+1}{k}}+O_{f,w,N}\left(\left(1+|\xi|\right)^{-\frac{N+1}{k}}\right),$
for some $a_{j}\in\mathbb{C}$. The above follows from [49, Proposition 3]. ∎
Proof of Corollary 4.
The proof follows immediately after applying Theorem 5 and Proposition 5.1 with $N=\lfloor\tfrac{\max\{k_{i}\}}{2}\rfloor$. ∎
Acknowledgements
A major discussion took place during the “Analytic Number Theory” program at the Institute Mittag-Leffler (IML) where we were in residence and we want to thank IML for extraordinary hospitality and work condition. SD thanks University of Michigan and Uppsala University, and SJ thanks University of York where significant parts of our work were completed. We want to thank Dmitry Kleinbock, Manuel Luethi, Andreas Strömbergsson, and Barak Weiss for helpful comments on an earlier draft of this paper. We thank Asaf Katz for giving us a proper reference regarding stationary phase. SD wants to thank Han Yu for many interesting discussion and Sam Chow for sending the Mathematica code from [12]. Finally, we want to thank our family for supporting us while writing the paper.